### v1

# Rigidity of proper quasi-homogeneous domains in positive flag manifolds

Abstract.
We show that, inside the Shilov boundary of any given Hermitian symmetric space of tube type, there is, up to isomorphism, only one proper domain whose action by its automorphism group is cocompact. This gives a classification of all closed proper manifolds locally modelled on such Shilov boundaries, and provides a positive answer, in the case of flag manifolds admitting a $\Theta$-positive structure, to a rigidity question of Limbeek and Zimmer.
1. Introduction
The general context of this paper is that of geometric structures. A manifold $M$ admits a $(G,X)$-structure, where $G$ is a real Lie group and $X$ is a $G$-homogeneous space, if there exists an atlas of charts on $M$ with values in $X$ whose changes of coordinates are given by elements of $G$. The manifold $M$ endowed with this structure is called a $(G,X)$-manifold. Geometric structures have been the object of deep work and fundamental theories since Klein’s Erlangen program, including Poincaré–Koebe’s uniformization theorem, Teichmüller theory, Thurston’s geometrization program and Perelman’s theorem; see [Gol10].
In this paper, we consider more specifically the case where the manifold $M$ is closed, the group $G$ is semisimple and the space $X$ is a flag manifold of $G$, i.e. a smooth projective variety on which $G$ acts transitively. There exist many examples of such $(G,X)$-manifolds $M$, including quotients of domains of discontinuity of Anosov representations [Fra05, GW12, KLP18]. However, it is in general difficult to classify all closed $(G,X)$-manifolds. To obtain classification results, one strategy is to assume, for instance, that the manifold $M$ is a quotient $\Omega/\Gamma$ of a “not too large” (i.e. proper) domain $\Omega$ of $X$ by a discrete subgroup $\Gamma$ of $G$ acting properly discontinuously and freely on $\Omega$. We say that the manifold $M$ is proper.
In the case where $G=\operatorname{PGL}(n,\mathbb{R})$ is the projective linear group and $X=\mathbb{P}(\mathbb{R}^{n})$ is the real projective space, this leads to the theory of divisible convex sets, whose main results and ideas are outlined in Section 1.1 below. This rich theory motivates general questions concerning “small” divisible open sets in arbitrary flag manifolds.
With this in mind, in this paper we investigate closed proper $(G,X)$-manifolds where $G$ is a Lie group of Hermitian type of tube type and $X$ is the Shilov boundary of the symmetric space of $G$ (in particular $X$ is a flag manifold of $G$); see Table 1. This reduces to the study of proper divisible domains of $X$. Our main result (Theorem 1.4) implies that these domains are actually subject to a strong rigidity, contrasting with the flexibility that can be observed in the real projective case.
1.1. Divisible convex sets in real projective geometry
Convex projective geometry generalizes real hyperbolic geometry. The objects of study are the open sets $\Omega\subset\mathbb{P}(\mathbb{R}^{n})$ that are properly convex, meaning they are bounded and convex in an affine chart. The group of elements of $\operatorname{PGL}(n,\mathbb{R})$ preserving $\Omega$ is called the automorphism group of $\Omega$ and is denoted by $\mathsf{Aut}(\Omega)$. The set $\Omega$ is divisible if there exists a discrete subgroup of $\mathsf{Aut}(\Omega)$ that acts cocompactly on $\Omega$, and quasi-homogeneous if $\mathsf{Aut}(\Omega)$ acts cocompactly on $\Omega$. Note that divisibility implies quasi-homogeneity.
Properly convex open subsets of the projective space that are divisible are called divisible convex sets and admit a rich theory, which was initiated in the 1960’s with work of Benzecri [Ben60]. This theory has since then been developed by numerous authors (see e.g. [Vin65, Gol90, CLT15]), in particular by Benoist in the early 2000’s [Ben00, Ben03, Ben05, Ben06]. The strictly convex case is well understood, while the nonstrictly convex case is still in expansion [Isl, CLM20, Zim23, Bla]. See [Ben08] for a survey on this theory and more references.
By [Vey70, Ben01], the theory of divisible convex sets reduces to the case where $\Omega$ is irreducible. In this case, either $\Omega$ is symmetric, i.e., there exists a symmetry in $\mathsf{Aut}(\Omega)$ at every point of $\Omega$, or $\mathsf{Aut}(\Omega)$ is a discrete Zariski-dense subgroup of $\operatorname{PGL}(n,\mathbb{R})$; see [Vin65, Koe99, Ben03].
There is a list of all irreducible symmetric domains in any dimension [Koe99]. All domains of this list identify with Riemannian symmetric spaces, and their isometry group and automorphism group (as properly convex open subsets in projective space) coincide. The simplest example is the real hyperbolic space $\mathbb{H}^{n-1}$ embedded in the projective space $\mathbb{P}(\mathbb{R}^{n})$ via the Klein model. In this case, we are in the framework of real hyperbolic geometry.
There exists cocompact lattices $\Gamma$ of $\operatorname{PO}(n-1,1)$ that admit nontrivial deformations into $\operatorname{PGL}(n,\mathbb{R})$ [JM87]. By Koszul’s openness Theorem [Kos68], the image of any small deformation of $\Gamma$ in $\operatorname{PGL}(n,\mathbb{R})$ still divides a divisible convex set, which can be nonsymmetric. There also exist several constructions of nonsymmetric irreducible divisible convex sets; the first examples were constructed by Kac–Vinberg in dimension 2, using Coxeter groups [VK67]. Kapovich later constructed examples that are not quasi-isometric to any symmetric space, and whose automorphism group is discrete and Gromov-hyperbolic, in any dimension $n-1\geq 4$ [Kap07]. Examples with a discrete and non-Gromov-hyperbolic automorphism group were constructed in projective dimension 3,4,5,6 by Benoist [Ben06] and 3 by Ballas–Danciger–Lee [BDL18], and recently, in any projective dimension $n-1\geq 3$ by Blayac–Viaggi [BV23].
Even though it has recently been proven that divisible convex sets in the projective space are subject to a certain rigidity [Zim23], the diversity of nonsymmetric examples highlights the importance of general results concerning them, beyond those on cocompact actions on Riemannian symmetric spaces.
1.2. Flag manifolds and rigidity
The real projective space is an example of a flag manifold, i.e. a compact homogeneous space $G/P$ where $G$ is a noncompact real semisimple linear Lie group and $P$ a parabolic subgroup of $G$. The question of whether the theory of divisible convex sets generalizes to other flag manifolds than the real projective space was asked by W. van Limbeek and A. Zimmer. There are natural notions of convexity, properness, divisibility and quasi-homogeneity for domains (i.e. connected open sets) in flag manifolds, generalizing those in the projective space (see Section 4.1.2 and [Zim18a]). By “natural”, we mean that the following property, already true in the projective space (by [Sho84]), remains true in general flag manifolds: proper quasi-homogeneous domains are convex.
One can first consider a particular class of flag manifolds, those that identify with a symmetric space of compact type, with isometry group a maximal compact subgroup of $G$. These flag manifolds are called Nagano spaces, extrinsinc symmetric spaces or symmetric $R$-spaces depending on the authors, and their list is known [Nag65]. There is a wealth of literature about these manifolds, starting with [KN64, Tak65, TK68, Mak73]. Nagano has classified symmetric spaces $\mathbb{X}$ of noncompact type that embed into their compact dual, which turn out to be Nagano spaces. Moreover, the image of $\mathbb{X}$ in its dual is a proper domain.
Nagano’s list provides examples of divisible convex sets in flag manifolds (beyond the projective space), but all these examples are symmetric. The issue is to determine if there exist nonsymmetric examples, as in the projective case. This is a general question, which applies to any flag manifold:
Question 1.1 ([LZ19]).
Given a noncompact real semisimple Lie group $G$, and a parabolic subgroup $P$ of $G$, are all proper divisible domain of $G/P$ symmetric?
One could also ask the same question, replacing “divisible” with “quasi-homogeneous”. A positive answer in the quasi-homogeneous case would imply a positive answer to Question 1.1.
First note that the study of proper quasi-homogeneous domains reduces to the case where $G$ is simple, by the following fact.
Fact 1.2 ([Zim18a, Thm 1.7]).
Let $G$ be a semisimple Lie group $G$ with trivial center and no compact factors, and write $G=G_{1}\times\cdots\times G_{k}$ where $G_{i}$ is a noncompact simple Lie group for all $1\leq i\leq k$. For any parabolic subgroup $P$ of $G$, there exist parabolic subgroups $P_{i}\leq G_{i}$ such that $P=P_{1}\times\cdots\times P_{k}$. Now let $\Omega\subset G/P$ be a proper quasi-homogeneous domain. Then there exist proper quasi-homogeneous domains $\Omega_{i}\subset G_{i}/P_{i}$ such that $\Omega=\Omega_{1}\times\cdots\times\Omega_{k}$.
Question 1.1 admits a negative answer in the case where $G=\operatorname{SO}(n,1)$ for $n\geq 3$, and $P$ is a minimal parabolic subgroup of $G$. In this case, the flag manifold $G/P$ is the conformal sphere and it admits proper divisible domains that are nonsymmetric (here “proper” simply means that the complement of the domain has nonempty interior). For instance, the limit set of a representation obtained by a deformation of the natural inclusion of a cocompact lattice in $\operatorname{SO}(2,1)$ into $\operatorname{SO}(3,1)$ (called quasi-Fuchsian) is a quasi-circle, and thus separates the two-sphere $G/P$ into two proper, divisible domains. The case where $G$ is the projective linear group $\operatorname{PGL}(n+1,\mathbb{R})$ and $G/P$ is the projective space $\mathbb{P}(\mathbb{R}^{n+1})$ also has a negative answer (see Section 1.1). However, in other flag manifolds $G/P$, one observes more rigidity, and Zimmer makes the conjecture that any divisible convex domain of $G/P$ is homogeneous [Zim18a, Conj. 2.6].
Question 1.1 has a complete answer for flag manifolds $G/P$ where $P$ is a nonmaximal proper parabolic subgroup:
Fact 1.3 ([Zim18a, Thm 1.5]).
Let $G$ be a noncompact real simple Lie group and $P\leq G$ a nonmaximal proper parabolic subgroup. Then there are no proper quasi-homogeneous domains in $G/P$.
Flag manifolds defined by maximal proper parabolic subgroups have also been studied: Question 1.1 has a positive answer for self-opposite Grassmannians $\operatorname{Gr}_{p}(\mathbb{R}^{2p})$ [LZ19] and for the complex projective space $\mathbb{P}(\mathbb{C}^{n})$ for $n\geq 3$ [Zim18b]. In both these cases, the authors show — under some additional assumption — that there is only one divisible proper domain in $G/P$ up to the action of $G$, and that this domain is symmetric.
In simultaneous joint work with A. Chalumeau [CG24], we give a positive answer to Question 1.1 when $G/P$ is the boundary of the pseudo-hyperbolic space $\mathbb{H}^{p,q}$ for $p\geq 2$ and $q\geq 1$ — also called the Einstein universe of signature $(p-1,q)$. The special case where $q=1$ admits a different proof than the general case, using causality in the Einstein universe of signature $(n-1,1)$. In the present paper, we generalize this causality approach to Shilov boundaries of Hermitian symmetric spaces of tube type (see the next section). These flag manifolds appear naturally in several contexts, such as Euclidean Jordan algebras and complex analysis (see e.g. [FK94]), or $\Theta$-positivity and higher Teichmüller Theory [GW16]. Here the proper parabolic subgroup $P$ is maximal, so Fact 1.3 does not apply. For any two transverse points $p,q\in G/P$, the open set consisting of all points of $G/P$ which are transverse to both $p$ and $q$ has several connected components, two of which are proper symmetric domains called diamonds (see Definition 4.6). The main result of the present paper (Theorem 1.4 below) is a positive answer to Question 1.1: we prove that in our context of Shilov boundaries there is, up to the action of $G$, only one proper quasi-homogeneous domain (hence only one proper divisible domain) in $G/P$, namely a diamond; in particular, it is symmetric.
Removing the properness assumption in Question 1.1 adds flexibility. Indeed, as mentioned above, there exist many examples of nonproper, nonsymmetric divisible domains of flag manifolds $G/P$, constructed e.g. as domains of discontinuity for Anosov representations [Fra05, GW12, KLP18].
1.3. Statement of the main theorem.
Let $G$ be a simple Lie group of Hermitian type of tube type and of real rank $r\geq 1$ (the complete list of the corresponding Lie algebras is given in Table 1). Let $\operatorname{\Delta}$ be set of simple restricted roots of $G$ and $\alpha_{r}\in\operatorname{\Delta}$ be the unique long root. Then the flag manifold $\mathscr{F}_{\{\alpha_{r}\}}=G/P_{\{\alpha_{r}\}}$ defined by $\{\alpha_{r}\}$ (see Section 2.2) is the Shilov boundary of the symmetric space $\mathbb{X}_{G}$ of $G$, denoted by $\mathbf{Sb}(G)$. Recall that the automorphism group $\mathsf{Aut}(\Omega)$ of a domain $\Omega\subset\mathbf{Sb}(G)$ is the group of all elements of $G$ that leave $\Omega$ invariant; we say that $\Omega$ is quasi-homogeneous if $\mathsf{Aut}(\Omega)$ acts cocompactly on $\Omega$.
Theorem 1.4.
Let $G$ be a simple Lie group of Hermitian type of tube type. Then any proper quasi-homogeneous domain of $\mathbf{Sb}(G)$ is a diamond.
As in [CG24, Sect. 6.3], the proof of Theorem 1.4 uses the causal structure of $\mathbf{Sb}(G)$, i.e. the existence (up to taking an index-two subgroup of $G$) of a $G$-equivariant smooth family $(c_{x})_{x\in\mathbf{Sb}(G)}$ of properly convex open cones in the tangent bundle $T(\mathbf{Sb}(G))$. Any proper domain $\Omega\subset\mathbf{Sb}(G)$ inherits a causal structure from the one of $\mathbf{Sb}(G)$. A generalization of Liouville’s classical theorem implies that $\mathsf{Aut}(\Omega)$ is commensurable to the conformal group of $\Omega$, that is, the group of diffeomorphisms $f:\Omega\rightarrow\Omega$ such that $d_{x}f(c_{x})=c_{f(x)}$ for all $x\in\Omega$ [Kan11]. Theorem 1.4 states that having a cocompact conformal group characterizes diamonds among proper domains of $\mathbf{Sb}(G)$.
More generally, if $G$ is a semisimple Lie group, Hermitian of tube type, with trivial center and no compact factors, then one can write $G=G_{1}\times\cdots\times G_{k}$, where $G_{i}$ is a noncompact simple Lie group, Hermitian of tube type, for all $1\leq i\leq k$. Then the Shilov boundary $\mathbf{Sb}(G)$ of $G$ is the flag manifold $\mathbf{Sb}(G_{1})\times\cdots\times\mathbf{Sb}(G_{k})$. By Fact 1.2 and Theorem 1.4, one directly gets:
Corollary 1.5.
Let $G$ be a semisimple Lie group of Hermitian type of tube type with trivial center and no compact factors, and write $G=G_{1}\times\cdots\times G_{k}$, where $G_{i}$ is a noncompact simple Lie group of Hermitian type of tube type for all $1\leq i\leq k$. Then for any proper quasi-homogeneous domain $\Omega\subset\mathbf{Sb}(G)$, there are diamonds $D_{i}\subset\mathbf{Sb}(G_{i})$ for $1\leq i\leq k$, such that
$\Omega=D_{1}\times\cdots\times D_{k}\subset\mathbf{Sb}(G_{1})\times\cdots\times\mathbf{Sb}(G_{k}).$
In particular, Question 1.1 has a positive answer for Shilov boundaries of Hermitian symmetric spaces of tube type.
Note that the flag manifolds studied in this paper are in the list of Nagano spaces, and a diamond of $\mathbf{Sb}(G)$ is a model for the noncompact dual of the compact symmetric space $\mathbf{Sb}(G)$.
1.4. $\Theta$-positive structures
Total positivity has been known and studied since the beginning of the 20th century for $\operatorname{SL}(N,\mathbb{R})$. It was generalized to real split semisimple Lie groups by Lusztig [Lus94]. On the other hand, isometry groups of irreducible Hermitian symmetric spaces of tube type were known to admit a causal structure.
Guichard–Wienhard generalized these two notions of total positivity and causality with their notion of $\Theta$-positive structure, where $\Theta$ is a subset of the simple roots of a semisimple Lie group $G$. They listed all pairs $(G,\Theta)$ such that $G$ admits a $\Theta$-positive structure [GW16, GW22]. The pairs $(G,\{\alpha_{r}\})$, where $G$ is a simple Lie group of Hermitian type of tube type and $\alpha_{r}$ is the simple root of $G$ defining the Shilov boundary $\mathbf{Sb}(G)$, constitute the only family of their list where the set $\Theta$ is a singleton (i.e. where the proper parabolic subgroup defined by $\Theta$ is maximal). Hence Theorem 1.4 and Fact 1.3 complete the classification of proper quasi-homogeneous domains in flag manifolds admitting a $\Theta$-positive structure:
Corollary 1.6.
Let $G$ be a real noncompact simple Lie group and $\Theta$ a subset of the simple restricted roots of $G$ such that $G$ admits a $\Theta$-positive structure. Then the following dichotomy holds:
  1. (1)
    If $|\Theta|=1$, then $G$ is Hermitian of tube type and $\mathscr{F}_{\Theta}=\mathbf{Sb}(G)$ admits exactly one proper quasi-homogeneous domain (up to the action of $G$), which is a diamond.
  2. (2)
    If $|\Theta|\geq 2$, then there does not exist any proper quasi-homogeneous domain in $\mathscr{F}_{\Theta}$.
Here we denote by $\mathscr{F}_{\Theta}$ the flag manifold defined by $(G,\Theta)$ (see Section 2.3). Again, by Fact 1.2, Question 1.1 has a positive answer for flag manifolds $G/P_{\Theta}$ with a $\Theta$-positive structure, where $G$ is a noncompact semisimple (not necessarily simple) Lie group and $\Theta$ a subset of the simple roots of $G$.
1.5. An application to $(G,G/P)$-structures.
In the real projective space $\mathbb{P}(\mathbb{R}^{n})$, a nonsymmetric properly convex open set $\Omega$ divided by a discrete subgroup $\Gamma$ of $\operatorname{PGL}(n,\mathbb{R})$ provides a closed $(\operatorname{PGL}(n,\mathbb{R}),\mathbb{P}(\mathbb{R}^{n}))$-manifold $\Omega/\Gamma$ that is not isomorphic to a compact quotient of the symmetric space of a noncompact Lie group.
The case of Shilov boundaries of irreducible Hermitian symmetric spaces of tube type is different. Any $(G,\mathbf{Sb}(G))$-manifold $M$ develops into an open subset of $\mathbf{Sb}(G)$; see Section 9. The manifold $M$ is proper if the image of its developing map is proper in $\mathbf{Sb}(G)$. We get the following corollary of Theorem 1.4, whose proof is given in Section 9:
Corollary 1.7.
Let $G$ be a simple Lie group of Hermitian type of tube type, and let $M$ be a closed connected $(G,\mathbf{Sb}(G))$-manifold. Assume moreover that $M$ is proper. Then the manifold $M$ identifies as a $(G,\mathbf{Sb}(G))$-manifold to a quotient $D/\Gamma$, where $D$ is a diamond of $\mathbf{Sb}(G)$ and $\Gamma$ is a cocompact lattice of $\mathsf{Aut}(D)$. Hence the manifold $M$ is a finite cover of
$\mathbb{S}^{1}\times(X_{L_{s}}/\Gamma^{\prime}),$
where $X_{L_{s}}$ is the symmetric space of the semisimple part $L_{s}$ of a Levi subgroup $L$ of $P_{\{\alpha_{r}\}}$, and $\Gamma^{\prime}$ is a cocompact lattice of $L_{s}$.
In particular, with the notation of Corollary 1.7, the developing map of $M$ is a diffeomorphism between the universal cover $\widetilde{M}$ of $M$ and $D$; we say that the manifold $M$ is Kleinian.
1.6. Comparison of two invariant metrics
In this paragraph we discuss an important step of the proof of Theorem 1.4 (see Proposition 1.9 below).
In the classical theory of divisible convex sets, the Hilbert metric is a powerful tool to understand the automorphism group of a properly convex domain $\Omega$, due to its $\mathsf{Aut}(\Omega)$-invariance. For a proper quasi-homogeneous domain $\Omega$, the Hilbert metric allows one to relate the geometry of $\partial\Omega$ to the dynamics of $\mathsf{Aut}(\Omega)$. Namely, in this case any extremal point is “visual” (in the sense of Section 7.1):
Fact 1.8 ([Vey70, Lem. 4]).
Let $\Omega$ be a properly convex domain of $\mathbb{P}(\mathbb{R}^{n})$. Assume that there exists a subgroup $H\leq\mathsf{Aut}(\Omega)$ such that $H$ acts cocompactly on $\Omega$. Then for any extremal point $p\in\partial\Omega$, there exists a sequence $(h_{n})\in H^{\mathbb{N}}$ such that for any compact subset $\mathscr{K}\subset\Omega$, the Hausdorff limit of $(h_{n}\cdot\mathscr{K})$ is $\{p\}$.
For every $z\in\mathbb{P}(\mathbb{R}^{n})$ we choose a lift $\widetilde{z}$ of $z$ in $\mathbb{R}^{n}$. Given two distinct points $x,y$ of $\mathbb{P}(\mathbb{R}^{n})$, we denote by $\ell_{x,y}$ the unique projective line through $x$ and $y$. Finally, given four points $a,b,c,d\in\mathbb{P}(\mathbb{R}^{n})$, we denote by $(a:x:y:b)$ their cross ratio. The Hilbert metric on a properly convex domain $\Omega\subset\mathbb{P}(\mathbb{R}^{n})$ can be defined in two ways: for any $x,y\in\Omega$, the distance $H_{\Omega}(x,y)$ is equal to both of the two quantities
  1. (1)
    $\inf\left\{\log(a:x:y:b)\mid a,b\in\Omega\cap\ell_{x,y},\ a,x,y,b\text{ aligned in this order}\right\}$;
  2. (2)
    $\sup_{\eta,\xi\in\Omega^{*}}\log\left|\frac{\xi(\widetilde{x})\eta(\widetilde{y})}{\xi(\widetilde{y})\eta(\widetilde{x})}\right|$, where $\Omega^{*}=\mathbb{P}\left(\{f\in\mathbb{R}^{n}\mid f(\widetilde{z})\neq 0\quad\forall z\in\overline{\Omega}\}\right)$.
One idea in this paper is to adapt some arguments and results from convex projective geometry, in particular Fact 1.8, to the context of a flag manifold $\mathscr{F}_{\Theta}$ different from the real projective space. This requires constructing an $\mathsf{Aut}(\Omega)$-invariant metric on a proper domain $\Omega\subset\mathscr{F}_{\Theta}$ that gives enough information about the structure of $\partial\Omega$. Definition (2) above of the Hilbert metric already admits generalizations to arbitrary flag manifolds $\mathscr{F}_{\Theta}$ [Zim18a], called Caratheodory metrics and defined from irreducible $\Theta$-proximal representations of $G$; see Section 6.2.1. These metrics turn out to be proper (i.e. closed balls are compact) as soon as $\Omega$ is dually convex (see Section 4.1.2 and the proof of [Zim18a, Thm 9.1]). However, in our case, these metrics do not provide enough information about $\partial\Omega$ to prove Theorem 1.4. Instead, we use the Kobayashi metric, which generalizes Definition (1) above and is denoted by $K_{\Omega}$ in this paper. Contrary to those of Caratheodory metrics, the definition of the Kobayashi metric requires more structure on $\mathscr{F}_{\Theta}$: namely, the existence of photons satisfying some conditions of invariance (Lemmas 5.4 and 5.5) and of abundance (Lemma 6.2). These properties are satisfied for Shilov boundaries of irreducible Hermitian symmetric spaces of tube type, but not in arbitrary flag manifolds.
Although we will not directly use Caratheodory metrics in the proof of Theorem 1.4, we will indirectly use them to deduce properties of the Kobayashi metric: with the notation of Section 2.4, we get the following proposition, whose proof is contained in Proposition 6.8 and Corollary 6.13:
Proposition 1.9.
Let $G$ be a simple Lie group of Hermitian type of tube type and let $(V,\rho)$ be a finite-dimensional real irreducible linear representation of $G$ with highest weight $\chi=N\omega_{\alpha_{r}}$, where $N\in\mathbb{N}_{>0}$ and $\omega_{\alpha_{r}}$ is the fundamental weight associated with the last simple root $\alpha_{r}$. Let $\Omega\subset\mathbf{Sb}(G)$ be a proper dually convex domain, and let $C_{\Omega}$ be the Caratheodory metric on $\Omega$ induced by $(V,\rho)$. Then one has
$K_{\Omega}\geq\frac{1}{N}C_{\Omega}.$
In particular, the metric $K_{\Omega}$ is proper.
Remark 1.10.
A Kobayashi metric can also be defined on proper domains of other flag manifolds, such as the Grassmannians [LZ19] and $\partial\mathbb{H}^{p,q}$ [CG24]. In this case, Proposition 1.9 still holds with a similar proof.
1.7. Outline of the proof of Theorem 1.4 and organization of the paper
We now describe the various sections of the paper in the order in which they appear, emphasizing on the key steps of the proof of Theorem 1.4. We denote by $G$ a simple Lie group of Hermitian type of tube type and of real rank $r\geq 1$, and by $\mathbf{Sb}(G)$ the Shilov boundary of the symmetric space $\mathbb{X}_{G}$ of $G$ (see Section 3).
1.7.1. Reminders on Lie groups and Shilov boundaries
In Section 2, we recall some facts on real Lie groups, flag manifolds, and $\Theta$-proximal representations. In Section 3, we recall and prove basic properties of Shilov boundaries of Hermitian symmetric spaces of tube type, and fix some notation for the paper.
1.7.2. Diamonds
In Section 4, following [Zim18a], we define the notions of properness and of automorphism groups for domains of $\mathbf{Sb}(G)$, in analogy with convex projective geometry. We introduce the previously mentioned diamonds (see Section 4.2). If $p$ lies in the standard affine chart $\mathbb{A}$ of $\mathbf{Sb}(G)$ (with the notation of Section 1.2), then causality allows us to define the future and the past of $p$ (see Section 3.3). For any point $q$ in the future of $p$, the intersection of the past of $q$ with the future of $p$ is a diamond, denoted by $\mathbf{D}(p,q)$; see Figure 2.
1.7.3. Photons.
In Section 5, we define a photon-generating action as an action of $\operatorname{SL}_{2}(\mathbb{R})$ on $\mathbf{Sb}(G)$ which is conjugate to the one induced by the $\mathfrak{sl}_{2}$-triple associated with the last simple root of $\mathfrak{g}$. A photon is then the unique closed orbit in $\mathbf{Sb}(G)$ of a photon-generating action. It is a topological circle whose properties are similar to those of projective lines in $\mathbb{P}(\mathbb{R}^{n})$ (Lemmas 5.4 and 5.5) and allow us to define and analyze the Kobayashi metric (see Section 6), and to understand the boundary of a proper domain in $\mathbf{Sb}(G)$ (see Section 7).
In Section 5.2 we introduce the intersection polynomials, to describe the intersection of a photon with a maximal proper Schubert subvariety of $\mathbf{Sb}(G)$, that is, a subset of $\mathbf{Sb}(G)$ of the form
(1.1) $\operatorname{Z}_{z}=\{x\in\mathbf{Sb}(G)\mid x\text{ is not transverse to }z\},$
where $z\in\mathbf{Sb}(G)$ (see Section 2.3). The complex roots of intersection polynomials are analyzed in Section 6.2.2, see Section 1.7.4 below.
1.7.4. The Kobayashi metric
In Section 6, we define the Kobayashi pseudo-metric on a domain $\Omega\subset\mathbf{Sb}(G)$, as mentioned above in Section 1.6, and investigate its basic properties (see Proposition 6.5). This pseudo-metric is a metric generating the standard topology whenever $\Omega$ is a proper domain of $\mathbf{Sb}(G)$ (Proposition 6.7).
A substantial part of Section 6 (more precisely, Sections 6.2.2 and 6.2.3) is devoted to proving Proposition 1.9. The complex roots of an intersection polynomial describe the intersection of complexifications of a photon and of a generic maximal proper Schubert subvariety of $\mathbf{Sb}(G)$; since this intersection is a singleton (Lemma 6.9), the intersection polynomials are split (Corollary 6.10). This observation allows us to compare two cross ratios, namely the one appearing in the definition of the Kobayashi metric and the other in that of the Caratheodory metric; see Lemma 6.11, from which follows Proposition 1.9.
1.7.5. A diamond containing $\Omega$.
Let $\Omega\subset\mathbf{Sb}(G)$ be a proper quasi-homogeneous domain. In Section 7, two families of points of $\partial\Omega$ are introduced: the visual points, defined by a dynamical property, and the $\mathscr{R}$-extremal points (see Definition 7.3), defined by a geometric one. The definition of $\mathscr{R}$-extremal points and that of the Kobayashi metric both involve photons, which allows us to prove an analogue of Fact 1.8, following the strategy of [LZ19]: namely, $\mathscr{R}$-extremal points are visual (Proposition 7.4).
By quasi-homogeneity, the domain $\Omega$ is dually convex [Zim18a, Thm 9.1], that is, for any $p\in\partial\Omega$ there is a supporting hypersurface to $\Omega$ at $p$ of the form $\operatorname{Z}_{z}$ (see (1.1)); see Section 4.1.2. Lemma 7.1 expresses a stronger geometric property for a visual point $p\in\partial\Omega$: this supporting hypersurface $\operatorname{Z}_{z}$ can be taken to be $\operatorname{Z}_{p}$ itself. This lemma, applied to two strongly $\mathscr{R}$-extremal points $p_{0},q_{0}\in\partial\Omega$ (see Definition 7.3 and Lemma 7.10), implies that $\Omega$ is contained in the diamond $\mathbf{D}(p_{0},q_{0})$ (Section 8).
1.7.6. Proving the equality.
In Section 8, with the notation of Section 1.7.5 above, we prove that $\Omega=\mathbf{D}(p_{0},q_{0})$. The key point is the inclusion $\mathsf{Aut}(\Omega)\leq\mathsf{Aut}(\mathbf{D}(p_{0},q_{0}))$, which implies, by quasi-homogeneity, that $\Omega$ is closed in (and hence equal to) the diamond $\mathbf{D}(p_{0},q_{0})$. This inclusion holds because any automorphism of $\Omega$ preserves the pair $\{p_{0},q_{0}\}$, which is proved in Proposition 8.2, and essentially characterizes $\mathbf{D}(p_{0},q_{0})$ (Fact 4.8).
1.7.7. An application to geometric structures
In Section 9, we prove Corollary 1.7, which is a consequence of Theorem 1.4.
Acknowledgements
I thank my PhD advisor Fanny Kassel for exposing this interesting topic to me and for her guidance, both in the proof and in the writing of this paper. I also thank Sami Douba for his valuable advice on the writing of this paper, and Jeff Danciger, Wouter van Limbeek and Andy Zimmer for interesting discussions on topics related to this work. I am grateful to Yves Benoist and François Labourie, the former for suggesting that I examine Question 1.1 for the $3$-dimensional Einstein universe, which, when generalized, led to [CG24] and the current paper, and the latter for sharing his insights on photons in positive flag manifolds. Finally, I warmly thank Adam Chalumeau for inspiring discussions in the context of our joint project [CG24].
2. Notation and basic reminders in Lie theory
2.1. Real and complex projective spaces, and cross ratio
In this section we set some notation for the elements of real and complex projective spaces, and we recall the definition of the cross ratio on $\mathbb{P}(\mathbb{R}^{2})$.
Given a finite-dimensional real vector space $V$, we will denote by $[v]$ the projection in $\mathbb{P}(V)$ of a vector $v\in V\smallsetminus\{0\}$. In the case where $V=\mathbb{R}^{2}$, we denote by $[t_{1}:t_{2}]$ the projection in $\mathbb{P}(\mathbb{R}^{2})$ of a vector $(t_{1},t_{2})\in\mathbb{R}^{2}\smallsetminus\{0\}$.
[…]
If $I\subset\mathbb{P}(\mathbb{R}^{2})$ is a proper open interval with (possibly equal) endpoints $t_{1}$ and $t_{2}$, then the Hilbert pseudo-metric on $I$ is denoted $H_{I}$ and defined as follows: for any pair $s_{1},s_{2}\in I$ such that $t_{1},s_{1},s_{2},t_{2}$ are aligned in this order (taking any order if $s_{1}=s_{2}$ or $t_{1}=t_{2}$), one has $H_{I}(s_{1},s_{2}):=\log(t_{1}:s_{1}:s_{2}:t_{2})$. If $I=\mathbb{P}(\mathbb{R}^{2})$, then $H_{I}$ is by convention the constant map equal to 00 on $I^{2}$.
In Section 6.2.2, we will also be led to consider the complex projective spaces. Since we will consider both real vector and their complexification, to avoid the confusion we denote by $\mathbb{P}_{c}(W)$ the complex projective space of a finite-dimensional complex vector space $W$, and by $[v]_{c}$ the projection in $\mathbb{P}_{c}(W)$ of a vector $v\in W\smallsetminus\{0\}$. We use the notation $[z_{1}:z_{2}]_{c}$ for the projection in $\mathbb{P}(\mathbb{C}^{2})$ of a vector $(z_{1},z_{2})\in\mathbb{C}^{2}\smallsetminus\{0\}$.
2.2. Reminders on Lie groups Lie algebras
In this section we recall some well-known facts about semisimple Lie groups and fix notations that will hold for the rest of the paper. We fix a real semisimple Lie gorup $G$ with Lie algebra $\mathfrak{g}$.
2.2.1. $\mathfrak{sl}_{2}$-triples
A triple $\mathsf{t}=(e,h,f)$ of elements of $\mathfrak{g}$ satisfying the equalities $[h,e]=2e$, $[h,f]=-2f$ and $[e,f]=h$ is called an $\mathfrak{sl}_{2}$-triple. There is a Lie algebras embedding $\operatorname{j}_{\mathsf{t}}:\mathfrak{sl}_{2}(\mathbb{R})\hookrightarrow\mathfrak{g}$ such that $\operatorname{j}_{\mathsf{t}}(\operatorname{E})=e$, $\operatorname{j}_{\mathsf{t}}(\operatorname{H})=h$ and $\operatorname{j}_{\mathsf{t}}(\operatorname{F})=f$, where
$\operatorname{E}=\begin{pmatrix}0&1\\ 0&0\end{pmatrix};\quad\operatorname{H}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix};\quad\operatorname{F}=\begin{pmatrix}0&0\\ -1&0\end{pmatrix}.$
2.2.2. Cartan decomposition
Let $B$ be the Killing form on $\mathfrak{g}$. Let $K\leq G$ be a maximal compact subgroup and $\mathfrak{p}$ be the $B$-orthogonal of the Lie algebra $\mathfrak{k}$ of $K$ in $\mathfrak{g}$. Then one has $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$. The Cartan involution of $\mathfrak{g}$ (with respect to $K$) is then the Lie algebra automorphism $\sigma:\mathfrak{g}\rightarrow\mathfrak{g}$ defined by $\sigma_{|\mathfrak{k}}=\operatorname{id}_{\mathfrak{k}}$ and $\sigma_{|\mathfrak{p}}=-\operatorname{id}_{\mathfrak{p}}$.
2.2.3. Restricted root system.
Let $\mathfrak{a}\subset\mathfrak{p}$ be a maximal abelian subspace, and $\mathfrak{g}_{0}$ the centralizer of $\mathfrak{a}$ in $\mathfrak{g}$. We denote by $\mathfrak{a}^{*}$ the space of all linear forms on $\mathfrak{a}$. For $\alpha\in\mathfrak{a}^{*}$, we define
$\mathfrak{g}_{\alpha}:=\{X\in\mathfrak{g}\mid[H,X]=\alpha(H)X\quad\forall H\in\mathfrak{a}\}.$
One has $[\mathfrak{g}_{\alpha},\mathfrak{g}_{\beta}]\subset\mathfrak{g}_{\alpha+\beta}$ for any $\alpha,\beta\in\mathfrak{a}^{*}$. If $\alpha\in\mathfrak{a}^{*}\smallsetminus\{0\}$ satisfies $\mathfrak{g}_{\alpha}\neq\{0\}$, then we say that $\alpha$ is a root of $(\mathfrak{g},\mathfrak{a})$. We denote by $\Sigma=\Sigma(\mathfrak{g},\mathfrak{a})$ the set of all the roots of $(\mathfrak{g},\mathfrak{a})$. One has $\mathfrak{g}=\mathfrak{g}_{0}\oplus\bigoplus_{\alpha\in\Sigma}\mathfrak{g}_{\alpha}$. We fix a fundamental system $\operatorname{\Delta}=\{\alpha_{1},\cdots,\alpha_{N}\}\subset\Sigma$, i.e. a family of roots such that any root of $\mathfrak{g}$ can be uniquely written as $\alpha=\sum_{i=1}^{N}n_{i}\alpha_{i}$, where the $n_{i}$ all have same sign for $1\leq i\leq N$. The elements of $\operatorname{\Delta}$ are called simple roots. The choice of a fundamental system determines a set of positive roots $\Sigma^{+}$, i.e. those roots $\alpha$ where the $n_{i}$ are all nonnegative.
For any $\alpha\in\Sigma$ and $X\in\mathfrak{g}_{\alpha}\smallsetminus\{0\}$, there exists a unique scalar multiple $X^{\prime}$ of $X$ such that $(X^{\prime},\sigma(-X^{\prime}),[\sigma(X^{\prime}),X^{\prime}])$ is an $\mathfrak{sl}_{2}$-triple. The element $[\sigma(X^{\prime}),X^{\prime}]$ does not depend on the choice of $X\in\mathfrak{g}_{\alpha}$, and is denoted by $h_{\alpha}$. The family $(h_{\alpha})_{\alpha\in\operatorname{\Delta}}$, forms a basis of $\mathfrak{a}$, whose dual basis in $\mathfrak{a}^{*}$ is denoted by $(\omega_{\alpha})_{\alpha\in\operatorname{\Delta}}$.
2.2.4. The restricted Weyl group
The restricted Weyl group $W$ of $G$ is the quotient $N_{K}(\mathfrak{a})/Z_{K}(\mathfrak{a})$ of the normalizer of $\mathfrak{a}$ in $K$ (for the adjoint action) by the centralizer of $\mathfrak{a}$ in $K$. It is a finite group generated by the $B$-orthogonal reflexions in $\mathfrak{a}$ with respect to the kernels of the simple roots. By duality with respect to $B$ (which induces a scalar product on $\mathfrak{a}$), the action of $W$ on $\mathfrak{a}$ induces an action on $\mathfrak{a}^{*}$ preserving $\Sigma$. There exists a unique $w_{0}\in W$, called the longest element, such that $w_{0}\cdot\Sigma^{+}=-\Sigma^{+}$. The element $i:\mathfrak{a}^{*}\rightarrow\mathfrak{a}^{*}$ defined as $i=-w_{0}$ is called the opposition involution, and satisfies $i(\operatorname{\Delta})=\operatorname{\Delta}$.
2.2.5. Parabolic subgroups.
Let $\Theta\subset\operatorname{\Delta}$ be a subset of the simple roots. The standard parabolic subgroup $P_{\Theta}$ (resp. the standard opposite parabolic subgroup $P_{\Theta}^{-}$) is defined as the normalizer in $G$ of the Lie algebra
(2.1) $\mathfrak{u}_{\Theta}^{+}:=\bigoplus_{\alpha\in\Sigma_{\Theta}^{+}}\mathfrak{g}_{\alpha}\quad\Bigg{(}\text{resp.\ }\mathfrak{u}_{\Theta}^{-}:=\bigoplus_{\alpha\in\Sigma_{\Theta}^{+}}\mathfrak{g}_{-\alpha}\Bigg{)},$
where $\Sigma_{\Theta}^{+}:=\Sigma^{+}\smallsetminus\operatorname{Span}(\operatorname{\Delta}\smallsetminus\Theta)$. By “standard”, we mean with respect to the above choices. One has
(2.2) $\operatorname{Lie}(P_{\Theta})=\mathfrak{g}_{0}\oplus\bigoplus_{\alpha\in\Sigma^{+}}\oplus\bigoplus_{\alpha\in\Sigma\smallsetminus\Sigma_{\Theta}^{+}}\mathfrak{g}_{-\alpha}.$
More generally, a parabolic subgroup of type $\Theta$ of $G$ is a conjugate of $P_{\Theta}$ in $G$.
The Levi subgroup associated with $\Theta$ is the reductive Lie group defined as the intersection $L_{\Theta}:=P_{\Theta}\cap P_{\Theta}^{-}$. The unipotent radical of $P_{\Theta}$ (resp. $P_{\Theta}^{-}$) is $U_{\Theta}^{+}:=\exp(\mathfrak{u}_{\Theta}^{+})$ (resp. $U_{\Theta}^{-}:=\exp(\mathfrak{u}_{\Theta}^{-})$). One then has $P_{\Theta}=U_{\Theta}^{+}\rtimes L_{\Theta}$ (resp. $P_{\Theta}^{-}=U_{\Theta}^{-}\rtimes L_{\Theta}$).
2.3. Flag manifolds
Let $G$ be a real semisimple Lie group. The flag manifold associated with $\Theta$ is the quotient space $\mathscr{F}_{\Theta}:=G/P_{\Theta}$. The flag manifold opposite to $\mathscr{F}_{\Theta}$ is $\mathscr{F}_{\Theta}^{-}:=G/P_{\Theta}^{-}$. The action of $G$ on $\mathscr{F}_{\Theta}\times\mathscr{F}_{\Theta}^{-}$ by left translations has exactly one open orbit $\mathscr{O}$, which is the orbit of $(P_{\Theta},P^{-}_{\Theta})$ and is dense. Two elements $x\in\mathscr{F}_{\Theta}$ and $y\in\mathscr{F}_{\Theta}^{-}$ are said to be transverse if $(x,y)\in\mathscr{O}$.
Given a point $y\in\mathscr{F}_{\Theta}^{-}$ (resp. $x\in\mathscr{F}_{\Theta}$), we let $\operatorname{Z}_{y}$ (resp. $\operatorname{Z}_{x}$) be the set of all elements of $\mathscr{F}_{\Theta}$ (resp. $\mathscr{F}_{\Theta}^{-}$) that are not transverse to $y$ (resp. to $x$). It defines an algebraic hypersurface of $\mathscr{F}_{\Theta}$ (resp. of $\mathscr{F}_{\Theta}^{-}$). The space $\mathscr{F}_{\Theta}\smallsetminus\operatorname{Z}_{y}$ is called an affine chart (or more classically a big Schubert cell) and is an open dense subset of $\mathscr{F}_{\Theta}$. The chart
(2.3) $\mathbb{A}:=\mathscr{F}_{\Theta}\smallsetminus\operatorname{Z}_{P_{\Theta}^{-}}$
is called the standard affine chart. The bijection
(2.4) $\varphi_{\mathsf{std}}:\begin{cases}\mathfrak{u}_{\Theta}^{-}&\overset{\sim}{\longrightarrow}\mathbb{A}\\ X&\longmapsto\exp(X)P_{\Theta}\end{cases}$
induces an affine structure on $\mathbb{A}$. Since $G$ acts transitively on $G/P^{-}_{\Theta}$, any affine chart $\mathscr{F}_{\Theta}\smallsetminus\operatorname{Z}_{y}$ with $y\in\mathscr{F}_{\Theta}^{-}$ admits an affine structure, which moreover depends only on $y$ (and not on the choice of $g\in G$ such that $y=gP_{\Theta}^{-}$).
If $i(\Theta)=\Theta$, then we say that $\mathscr{F}_{\Theta}$ is self-opposite. For any representative $k_{0}\in N_{K}(\mathfrak{a})$ of the longest element $w_{0}$, one has $P_{\Theta}=k_{0}P_{\Theta}^{-}k_{0}^{-1}$, and one can identify $\mathscr{F}_{\Theta}$ and $\mathscr{F}_{\Theta}^{-}$ via the $G$-equivariant diffeomorphism $\mathscr{F}_{\Theta}\rightarrow\mathscr{F}_{\Theta}^{-};\ hP_{\Theta}\mapsto hk_{0}P_{\Theta}^{-}$. We will always make this identification in this paper.
2.3.1. The automorphism group.
The group of all the Lie algebra automorphisms of $\mathfrak{g}$ is called the automorphism group of $\mathfrak{g}$ and denoted by $\mathsf{Aut}(\mathfrak{g})$. It is a Lie group with Lie algebra $\mathfrak{g}$. When $G$ is simple, the map $\operatorname{Ad}:G\rightarrow\mathsf{Aut}(\mathfrak{g})$ has finite kernel.
In general, the group $\mathsf{Aut}(\mathfrak{g})$ does not act on $\mathscr{F}_{\Theta}$. However, it admits a finite-index subgroup that does: indeed, any $g\in\mathsf{Aut}(\mathfrak{g})$ induces an automorphism $\psi_{g}$ of the fundamental system $\Delta$. This defines a group homomorphism $\mathsf{Aut}(\mathfrak{g})\rightarrow\mathsf{Aut}(\Delta)$, whose kernel is denoted by $\mathsf{Aut}_{1}(\mathfrak{g})$ and contains $\operatorname{Ad}(G)$. This kernel acts on $\mathscr{F}_{\Theta}$ for any $\Theta\subset\Delta$. This action extends the one of $G$, in the sense that for any $g\in G$ and for any $x\in\mathscr{F}_{\Theta}$, one has $g\cdot x=\operatorname{Ad}(g)\cdot x$. In particular $\ker(\operatorname{Ad})$ acts trivially on $\mathscr{F}_{\Theta}$.
2.4. Irreducible representations of semisimple Lie groups.
Let $(V,\rho)$ be a finite-dimensional real linear representation of a real semisimple Lie group $G$. We will denote by $\rho_{*}:\mathfrak{g}\rightarrow\operatorname{End}(V)$ the differential of $\rho$ at $\operatorname{id}$.
2.4.1. Restricted weights
For any $\lambda\in\mathfrak{a}^{*}$, we define
$V^{\lambda}:=\{v\in V\mid\rho(h)\cdot v=\lambda(h)v\quad\forall h\in\mathfrak{a}\}.$
If $V^{\lambda}\neq\{0\}$ we say that $\lambda$ is a restricted weight of $(V,\rho)$. Given $\alpha,\lambda\in\mathfrak{a}^{*}$, one has $\rho_{*}(X)\cdot V^{\lambda}\subset V^{\lambda+\alpha}$ for all $X\in\mathfrak{g}_{\alpha}$. For each $\alpha\in\operatorname{\Delta}$, the element $\omega_{\alpha}\in\mathfrak{a}^{*}$ introduced in Section 2.2.3 is called the fundamental weight associated with $\alpha$. The cone generated by the simple roots determines a partial ordering on $\mathfrak{a}^{*}$ given by
$\lambda\leq\lambda^{\prime}\Longleftrightarrow\lambda^{\prime}-\lambda\in\sum_{\alpha\in\operatorname{\Delta}}\mathbb{R}_{+}\alpha.$
If $(V,\rho)$ is a finite-dimensional real irreducible linear representation of $G$, then the set of restricted weights of $(V,\rho)$ admits a unique maximal element for that ordering (see [GW09, Cor. 3.2.3]). This element is called the highest weight of $\rho$, and denoted by $\chi_{\rho}$ or $\chi$.
2.4.2. Proximality and $\Theta$-proximal representations.
An automorphism $g\in\operatorname{GL}(V)$ is said to be proximal in $\mathbb{P}(V)$ if it has a unique eigenvalue of maximal modulus and if the corresponding eigenspace is one-dimensional.
Let $\Theta\subset\operatorname{\Delta}$ be a nonempty subset of the simple roots. We say that $(V,\rho)$ is $\Theta$-proximal if $\rho(G)$ contains a proximal automorphism of $V$ and $\{\alpha\in\operatorname{\Delta}\mid\langle\chi,\alpha\rangle>0\}=\Theta$ (see [GGKW17]). In this case, we denote by $V^{\chi}$ the weight space associated with $\chi$ (it is automatically one-dimensional), and $V^{<\chi}$ the sum of all other weight spaces.
Proposition 2.1 ([GGKW17, Prop. 3.3]).
One has:
  1. (1)
    The stabilizer of $V^{\chi}$ in $G$ (resp. $V^{<\chi}$) is $P_{\Theta}$ (resp. $P_{\Theta}^{-}$).
  2. (2)
    The maps $g\mapsto\rho(g)\cdot V^{\chi}$ and $g\mapsto\rho(g)\cdot V^{<\chi}$ induce two $\rho$-equivariant maps:
    $\iota:\mathscr{F}_{\Theta}\longrightarrow\mathbb{P}(V)\text{ and }\iota^{*}:\mathscr{F}_{\Theta}^{-}\longrightarrow\mathbb{P}(V^{*}).$
    Two elements $(x,\xi)\in\mathscr{F}_{\Theta}\times\mathscr{F}_{\Theta}^{-}$ are transverse if, and only if, their images $\iota(x)$ and $\iota^{*}(\xi)$ are.
Note that for $\nu\in V\smallsetminus\{0\}$ and $f\in V^{*}\smallsetminus\{0\}$, the transversality of $[\nu]\in\mathbb{P}(V)$ and $[f]\in\mathbb{P}(V^{*})$ is equivalent to $f(\nu)\neq 0$.
We will identify the space $\mathbb{P}(V^{*})$ with the set $\mathbb{P}^{\prime}(V)$ of projective hyperplanes of $\mathbb{P}(V)$, via the bijection $\mathbb{P}(V^{*})\overset{\sim}{\rightarrow}\mathbb{P}^{\prime}(V)$ defined by $[f]\mapsto\mathbb{P}(\ker(f)).$ With this identification and the notations of Proposition 2.1, two elements $(x,\xi)\in\mathscr{F}_{\Theta}\times\mathscr{F}_{\Theta}^{-}$ are transverse if, and only if, one has $\iota(x)\notin\iota^{*}(\xi)$.
3. Shilov boundaries and causality
If $G$ is a simple Lie group of Hermitian type of tube type, that is, if the symmetric space $\mathbb{X}_{G}$ of $G$ is irreducible and Hermitian of tube type, then we will say that $G$ is a $\operatorname{HTT}$ Lie group, and $\mathfrak{g}$ a $\operatorname{HTT}$ Lie algebra. In this section, we fix an $\operatorname{HTT}$ Lie group $G$ with Lie algebra $\mathfrak{g}$ and prove useful preliminary results.
3.1. Strongly orthogonal roots and root system
Two roots $\alpha,\beta\in\Sigma$ are called strongly orthogonal if neither $\alpha+\beta$ nor $\alpha-\beta$ is a root. Since $G$ is of tube type, there exists a (maximal) set $\{2\varepsilon_{1},\cdots,2\varepsilon_{r}\}\subset\Sigma$ of strongly orthogonal roots, such that the set $\operatorname{\Delta}=\{\alpha_{1},\cdots,\alpha_{r}\}$ is a fundamental system of $\Sigma$, where $\alpha_{i}=\varepsilon_{i}-\varepsilon_{i+1}$ for $i<r$ and $\alpha_{r}=2\varepsilon_{r}$. The system $\Sigma$ is then of type $C_{r}$ (see e.g. [FK94]):
$\Sigma=\{\pm\varepsilon_{i}\pm\varepsilon_{j}\mid 1\leq i\leq j\leq r\};\quad\Sigma^{+}=\{\varepsilon_{i}\pm\varepsilon_{j}\mid 1\leq i<j\leq r\}\cup\{2\varepsilon_{i}\mid 1\leq i\leq r\}.$
This is actually a characterization of $\operatorname{HTT}$ Lie groups:
Fact 3.1.
A simple Lie group $G$ is $\operatorname{HTT}$ if, and only if, its root system is of type $C_{r}$.
Recall the notation of Section 3.1. Then:
Lemma 3.2.
Let $G$ be an $\operatorname{HTT}$ Lie group. Let $\beta\in\Sigma_{\{\alpha_{r}\}}^{+}\smallsetminus\{\alpha_{r}\}$, and let $Y_{\beta}\in\mathfrak{g}_{\beta}$. Then there exists $Z\in\operatorname{Lie}\left(P_{\{\alpha_{r}\}}\right)$ such that $[v^{-},Z]=0$ and $\operatorname{Ad}(\exp\left(Y_{\beta}\right))v^{-}=v^{-}+Z$.
  • Proof.
    One has
    $\operatorname{Ad}\left(\exp\left(Y_{\beta}\right)\right)v^{-}=\exp\left(\operatorname{ad}(Y_{\beta})\right)v^{-}=\sum_{k=0}^{\infty}\frac{\operatorname{ad}(Y_{\beta})^{k}v^{-}}{k!}.$
    We compute the terms $\operatorname{ad}(Y_{\beta})^{k}v^{-}$ for $k\in\mathbb{N}$. Recall that $\Sigma$ is of type $C_{r}$. In the notation of Section 2.2, there are thus several possible values for $\beta$:
    If $\beta=2\varepsilon_{i}$ for some $11\leq i\leq r-$ 1 or $\varepsilon_{i}+\varepsilon_{j}$ for some $11\leq i<j\leq r-$ 1, then $\beta-\alpha_{r}\notin\Sigma$, so $[Y_{\beta},v^{-}]\in\mathfrak{g}_{\beta-\alpha_{r}}=\{0\}$. Hence for any $k\geq 1$, one has $\operatorname{ad}(Y_{\beta})^{k}v^{-}=0$, and so $\operatorname{Ad}(e^{Y_{\beta}})v^{-}=v^{-}$. In this case, we have $Z=0$.
If $\beta=\varepsilon_{i}+\varepsilon_{r}$ for some $1\leq i<r$, then $3\beta-\alpha_{r}=3\varepsilon_{i}+\varepsilon_{r}\notin\Sigma$. Hence for any $k\geq 3$ one has $\operatorname{ad}(Y_{\beta})^{k}v^{-}=0$. From this we deduce that $e^{\operatorname{ad}(Y_{\beta})}v^{-}=v^{-}+Z$, with
$Z:=[Y_{\beta},v^{-}]+\frac{1}{2}\left[Y_{\beta},[Y_{\beta},v^{-}]\right]\in\mathfrak{g}_{\beta-\alpha_{r}}\oplus\mathfrak{g}_{2\beta-\alpha_{r}}=\mathfrak{g}_{\varepsilon_{i}-\varepsilon_{r}}\oplus\mathfrak{g}_{2\varepsilon_{i}}\subset\operatorname{Lie}\left(P_{\{\alpha_{r}\}}\right)$
[…]
If $\Theta=\{\alpha_{r}\}$, then the flag manifold $\mathscr{F}_{\Theta}$ is called the Shilov boundary of $\mathbb{X}_{G}$ and we will denote it by $\mathbf{Sb}(G)$.
Since $\operatorname{\Delta}$ is of type $C_{r}$, its automorphism group is trivial (see e.g. [Kna96]). Hence the opposition involution is trivial and the flag manifold $\mathbf{Sb}(G)$ is self-opposite. Moreover, the groups $\mathsf{Aut}_{1}(\mathfrak{g})$ and $\mathsf{Aut}(\mathfrak{g})$ coincide, hence the group $\mathsf{Aut}(\mathfrak{g})$ acts on $\mathscr{F}_{\{\alpha_{r}\}}$.
3.1.1. Notation.
When $G$ is an $\operatorname{HTT}$ Lie group and $\Theta=\{\alpha_{r}\}$, we will always use the following simplified notation: $\mathfrak{u}^{\pm}=\mathfrak{u}_{\{\alpha_{r}\}}^{\pm}$, $\mathfrak{l}=\mathfrak{l}_{\{\alpha_{r}\}}$, $U^{\pm}=U_{\{\alpha_{r}\}}^{\pm}$, $P=P_{\Theta}$, $L=L_{\Theta}$. Note that the Lie algebras $\mathfrak{u}^{\pm}$ are abelian.
Since $P$ is a maximal proper parabolic subgroup of $G$, the center of $L$ is at most one-dimensional, and one can write $L=\mathbb{R}\times L_{s}$, where $L_{s}$ is the semisimple part of $L$. The possible values of the Lie algebra $\mathfrak{l}_{s}$ of $L_{s}$ are listed in Table 1.
3.2. Dilations and translations.
Let $G$ be an $\operatorname{HTT}$ Lie group. There exists $H_{0}$ in the center of $\mathfrak{l}$ such that $\mathfrak{u}^{\pm}$ is the root space of $\operatorname{ad}(H_{0})$ for the eigenvalue ±$\pm 1$ (see e.g. [Kan98]). For all $t\in\mathbb{R}_{>0}$ we define $\ell_{0}(t)=\exp\left(-\log\left(\sqrt{t}\right)H_{0}\right)\in L$. The map $\ell_{0}$ continuously extends to $\operatorname{id}$ at 00. The element $\operatorname{Ad}(\ell_{0}(t))$ acts on $\mathfrak{u}^{\pm}$ by
(3.1) $\operatorname{Ad}(\ell_{0}(t))X=\begin{cases}tX\quad&\forall X\in\mathfrak{u}^{-};\\ \frac{1}{t}X\quad&\forall X\in\mathfrak{u}^{+}.\end{cases}$
Hence any positive dilation of $\mathbb{A}$ (see (2.3)) at $P=\varphi_{\mathsf{std}}(0)$ can be realized as the restriction to $\mathbb{A}$ of a map of the form $x\mapsto\ell_{0}(t)\cdot x$ of $\mathbf{Sb}(G)$ for some $t\in\mathbb{R}_{>0}$.
Moreover, since $\mathfrak{u}^{-}$ is abelian, any translation in $\mathbb{A}$ is realized as left multiplication by an element of $U^{-}\leq G$. Each time we will talk about a translation in $\mathbb{A}$, it will mean that we apply a multiplication by an element of $U^{-}$.
According to the two previous paragraphs, for any dilation $d$ at a point $x_{0}\in\mathbb{A}$, there exists $g\in G$ such that $d$ coincides with the restriction of the map $x\mapsto g\cdot x$ of $\mathbf{Sb}(G)$ to $\mathbb{A}$. Each time we will talk about dilating at $x_{0}$ in $\mathbb{A}$, it will mean that we apply such a map.
Remark 3.3.
It is not true for a general simple Lie group $G$ and subset of the simple roots $\Theta$ that there exists $H_{0}\in\mathfrak{l}_{\Theta}$ such that $\operatorname{ad}(H_{0})X^{\pm}=\pm X^{\pm}$ for all $X^{\pm}\in\mathfrak{u}_{\Theta}^{\pm}$. This property is equivalent to $\mathfrak{u}^{\pm}_{\Theta}$ being abelian. When this is the case, the algebra $\mathfrak{g}$ admits a decomposition $\mathfrak{g}=\mathfrak{g}_{-1}\oplus\mathfrak{g}_{0}\oplus\mathfrak{g}_{1}$, with $\mathfrak{g}_{-1}=\mathfrak{u}_{\Theta}^{-}$, $\mathfrak{g}_{0}=\mathfrak{l}_{\Theta}$ and $\mathfrak{g}_{1}=\mathfrak{u}_{\Theta}^{+}$, and $[\mathfrak{g}_{k},\mathfrak{g}_{k^{\prime}}]\subset\mathfrak{g}_{k+k^{\prime}}$ for $k,k^{\prime}\in\{-1,0,1\}$, with $\mathfrak{g}_{m}:=\{0\}$ if $m\notin\{-1,0,1\}$. The element $H_{0}$ is then called the characteristic element of the graded Lie algebra $\mathfrak{g}$, that is, each space $\mathfrak{g}_{k}$ with $k\in\{-1,0,1\}$ is the eigenspace of $\operatorname{ad}(H_{0})$ for the eigenvalue $k$.
3.3. An invariant cone and causality
The identity component $L^{0}$ of $L$ acts irreducibly on $\mathfrak{u}^{-}$. By [Ben00, Prop. 4.7] applied to this action, there exists an open $L^{0}$-invariant properly convex cone $c^{0}$ in $\mathfrak{u}^{-}$. This cone is defined as the interior of the convex hull in $\mathfrak{u}^{-}$ of the orbit $\operatorname{Ad}(L^{0})\cdot v^{-}$. Note that the existence of such a cone endows $\mathbf{Sb}(G)$ with an invariant causal structure (see [Kan06]).
For any $x\in\mathbb{A}$ (recall (2.3))), there exists a unique $X\in\mathfrak{u}^{-}$ such that $x=\exp(X)P$. We define
$\mathbf{I}^{+}(x):=\varphi_{\mathsf{std}}(X+c^{0})=\exp(c^{0})\cdot x\text{, the \emph{future} of }x;$
$\mathbf{I}^{-}(x):=\varphi_{\mathsf{std}}(X-c^{0})=\exp(-c^{0})\cdot x\text{, the \emph{past} of }x;$
$\mathbf{J}^{+}(x):=\varphi_{\mathsf{std}}(X+\overline{c^{0}})=\exp(\overline{c^{0}})\cdot x\text{, the \emph{large future} of }x;$
$\mathbf{J}^{-}(x):=\varphi_{\mathsf{std}}(X-\overline{c^{0}})=\exp(-\overline{c^{0}})\cdot x\text{, the \emph{large past} of }x;$
$\mathbf{C}^{+}(x):=\varphi_{\mathsf{std}}(X+\partial c^{0})=\exp(\partial c^{0})\cdot x\text{, the \emph{future lightcone} of }x;$
$\mathbf{C}^{+}(x):=\varphi_{\mathsf{std}}(X-\partial c^{0})=\exp(-\partial c^{0})\cdot x\text{, the \emph{past lightcone} of }x;$
$\mathbf{C}(x):=\mathbf{C}^{+}(x)\cup\mathbf{C}^{-}(x)\text{, the \emph{lightcone} of }x.$
The following fact is well-known:
Fact 3.4.
The lightcone $\mathbf{C}(x)$ of $x\in\mathbb{A}$ is always contained in $\operatorname{Z}_{x}\cap\mathbb{A}$, and $\mathbf{I}^{\pm}(x)$ are connected components of $\mathbb{A}\smallsetminus\operatorname{Z}_{x}$.
The past, the future and the lightcone of a point $x\in\mathbb{A}$ are not invariant under the stabilizer $\operatorname{Stab}_{G}(x)$ of $x$ in $G$. However, they are locally invariant:
Lemma 3.5.
Let $x\in\mathbb{A}$ and $g\in G$ be such that $g\cdot P\in\mathbb{A}$. Then for any $\delta_{1}\in\{-,+\}$, there exist $\delta_{2}\in\{-,+\}$ and a neighborhood $\mathscr{U}$ of $x$ such that $g\cdot(\mathscr{U}\cap\mathbf{I}^{\delta_{1}}(x))\subset\mathbf{I}^{\delta_{2}}(g\cdot x)$.
  • Proof.
    Noticing that $\exp(X)\cdot\mathbf{I}^{\delta_{1}}(P)=\mathbf{I}^{\delta_{1}}(\exp(X)P)$ for all $X\in\mathfrak{u}^{-}$, we may assume that $x=P$. Let us prove the lemma for $\delta_{1}=+$, the proof being the same for $\delta_{1}=-$.
    Since $gP\in\mathbb{A}$, by (2.4) one can write $g=g^{\prime}\exp(Y)$, with $Y\in\mathfrak{u}^{+}$ and $g^{\prime}\in P^{-}$. There exists a neighborhood $\mathscr{U}$ of $P$, convex in $\mathbb{A}$, such that $g\cdot\mathscr{U}\subset\mathbb{A}$. Hence we have $\exp(Y)\cdot\mathscr{U}\subset(g^{\prime})^{-1}\cdot\mathbb{A}=\mathbb{A}$. Recall the map $\ell_{0}:\mathbb{R}_{\geq 0}\rightarrow L$ defined in Section 3.2. Since $\mathscr{U}$ is convex, by (3.1), one has $\ell_{0}(t)\cdot\mathscr{U}\subset\mathscr{U}$ for all $t\in[0,1]$. Then:
    (3.2) $\exp(tY)\cdot\mathscr{U}=\ell_{0}(t)^{-1}\exp(Y)\ell_{0}(t)\cdot\mathscr{U}\subset\ell_{0}(t)^{-1}\exp(Y)\cdot\mathscr{U}\subset\ell_{0}(t)^{-1}\cdot\mathbb{A}\subset\mathbb{A}.$
    Since $\mathscr{U}$ and $\mathbf{I}^{+}(P)$ are both convex, the set $\mathscr{U}\cap\mathbf{I}^{+}(P)$ is connected. For this reason, by (3.2) and since $U^{+}$ stabilizes $\operatorname{Z}_{P}$, the set $\exp(tY)\cdot(\mathscr{U}\cap\mathbf{I}^{+}(P))$ is contained in a connected component of $\mathbb{A}\smallsetminus\operatorname{Z}_{P}$ for all $t\in[0,1]$, let us denote this component by $\mathscr{V}$. By continuity, this component $\mathscr{V}$ does not depend on $t$. In particular, for $t=0$, one has $\exp(tY)=\operatorname{id}$, so $\mathscr{V}=\mathbf{I}^{+}(P)$. Hence $\exp(tY)\cdot(\mathscr{U}\cap\mathbf{I}^{+}(P))\subset\mathbf{I}^{+}(P)$.
    Now, since $g^{\prime}\in P^{-}$, there exist $X\in\mathfrak{u}^{-}$ and $\ell\in L$, such that $g^{\prime}=\exp(X)\ell$. But the element $\ell\in L$ either preserves $\mathbf{I}^{+}(P)$ or maps it to $\mathbf{I}^{-}(P)$ (see e.g. [GW22, Cor. 5.3]). On the other hand, since $U^{-}$ is abelian, one has $\exp(X)\cdot\mathbf{I}^{\pm}(P)=\mathbf{I}^{+}(\exp(X)P)$. Then:
    $g\cdot(\mathscr{U}\cap\mathbf{I}^{+}(P))$ $\subset g^{\prime}\cdot\mathbf{I}^{+}(P)$
    $=\exp(X)\ell\cdot\mathbf{I}^{+}(P)$
    =$=\begin{cases}\mathbf{I}^{+}(\exp(X)P)=\mathbf{I}^{+}(gP)\quad\text{if }\ell\cdot\mathbf{I}^{+}(P)=\mathbf{I}^{+}(P);\\ \mathbf{I}^{-}(\exp(X)P)=\mathbf{I}^{-}(gP)\quad\text{if }\ell\cdot\mathbf{I}^{+}(P)=\mathbf{I}^{-}(P).\end{cases}\hfill\square$
The future and the past of a point satisfy the following causality properties:
Fact 3.6.
For all $x,y,z\in\mathbb{A}$, one has:
  1. *
    (reflexivity) $x\in\mathbf{J}^{+}(y)\Leftrightarrow y\in\mathbf{J}^{-}(x)$;
  2. *
    (antisymmetry) $[x\in\mathbf{J}^{+}(y)\text{ and }y\in\mathbf{J}^{+}(x)]\Rightarrow x=y$;
  3. *
    (transitivity) $[x\in\mathbf{J}^{\pm}(y)\text{ and }y\in\mathbf{J}^{\pm}(z)\text{ }]\Rightarrow x\in\mathbf{J}^{\pm}(z)$
These properties are also true replacing “$\mathbf{J}$” with “$\mathbf{C}$”. Reflexivity and transitivity are also true replacing “$\mathbf{J}$” with “$\mathbf{I}$”.
3.4. Examples
The complete list of Shilov boundaries associated with $\operatorname{HTT}$ Lie algebras is given in Table 1.
$\mathfrak{g}$ $\mathbf{Sb}(G)$ $\mathfrak{l}_{s}$
$\mathfrak{so}(2,n)$, $n\geq 3$ $\operatorname{Ein}^{n-1,1}$ $\mathfrak{so}(n-1,1)$
$\mathfrak{sp}(2r,\mathbb{R})$ $\operatorname{Lag}_{r}(\mathbb{R}^{2r})$ $\mathfrak{sl}(r,\mathbb{R})$
$\mathfrak{u}(r,r)$ $\operatorname{Lag}_{r}(\mathbb{C}^{2r})$ $\mathfrak{sl}(r,\mathbb{C})$
$\mathfrak{so}^{*}(4r)$ $\operatorname{Lag}_{r}(\mathbb{H}^{2r})$ $\mathfrak{sl}(r,\mathbb{H})$
$\mathfrak{e}_{7(-25)}$ $\operatorname{Lag}_{3}(\mathbb{O}^{6})$ $\mathfrak{e}_{6(-26)}$
Table 1. Shilov boundaries associated with all the $\operatorname{HTT}$ Lie algebras.
Let us explain the notations in the table. For the notation $\mathfrak{e}_{7(-25)}$ and $\mathfrak{e}_{6(-26)}$, see [FK94].
3.4.1. The Lagrangians.
Let $\mathbb{K}=\mathbb{R},\mathbb{C}$ or $\mathbb{H}$, and let $r\geq 2$. We will define the set of Lagrangians as a submanifold of the space $\operatorname{Gr}_{r}(\mathbb{K}^{2r})$ of all the $r$-dimensional vector spaces of $\mathbb{K}^{2r}$. Let
$J_{\mathbb{K}}=\begin{pmatrix}0&-I_{r}\\ I_{r}&0\end{pmatrix}\in\operatorname{GL}(2r,\mathbb{K}),$
where $I_{r}$ is the identity matrix of size $r$. Given a matrix $g\in M_{2r}(\mathbb{K})$, we denote by $\overline{g}$ the matrix whose $(i,j)$-th entry is the conjugate (in $\mathbb{K}$) of the $(i,j)$-th entry of $g$. Let $G_{\mathbb{K}}:=\left\{g\in\operatorname{SL}(2r,\mathbb{K})\mid\ ^{t}\overline{g}J_{\mathbb{K}}g=J_{\mathbb{K}}\right\}$. One has $G_{\mathbb{R}}=\operatorname{Sp}(2r,\mathbb{R})$, $G_{\mathbb{C}}=\operatorname{SU}(r,r)$ and $G_{\mathbb{H}}=\operatorname{SO}^{*}(2r)$.
Let $\mathbf{b}$ be the bilinear form whose matrix in the canonical basis of $\mathbb{K}^{2r}$ is $J_{\mathbb{K}}$. The space $\operatorname{Lag}_{r}(\mathbb{K}^{2r})$ is the space of Lagrangians of $(\mathbb{K},\mathbf{b})$, i.e. the space of totally isotropic $r$-planes of $\mathbb{K}^{2r}$. It is a model for $\mathbf{Sb}(G_{\mathbb{K}})$.
If $(e_{1},\cdots e_{2r})$ is the canonical basis of $\mathbb{K}^{2r}$, then $\mathbf{Sb}(G_{\mathbb{K}})=G_{\mathbb{K}}/P$, where $P$ is the stabilizer in $G_{\mathbb{K}}$ of $\xi_{0}:=\operatorname{Span}(e_{1},\cdots,e_{r})$, and $P^{-}$ is the stabilizer in $G_{\mathbb{K}}$ of $\xi_{\infty}:=\operatorname{Span}(e_{r+1},\cdots,e_{2r})$. This model gives the following descriptions of $\mathfrak{u}^{\pm}$ and $L$:
$\mathfrak{u}^{-}$ =$=\left\{\begin{pmatrix}0_{r}&0\\ X&0_{r}\end{pmatrix}\mid X\in\operatorname{Mat}_{r}(\mathbb{K}),\ ^{t}\overline{X}=X\right\};$
$\mathfrak{u}^{+}$ =$=\left\{\begin{pmatrix}0_{r}&X\\ 0&0_{r}\end{pmatrix}\mid X\in\operatorname{Mat}_{r}(\mathbb{K}),\ ^{t}\overline{X}=X\right\};$
$L$ =$=\big{\{}\operatorname{diag}(A,\ ^{t}\overline{A}^{-1})\mid A\in\operatorname{GL}(r,\mathbb{K})\big{\}}.$
Hence the affine chart $\mathbb{A}=\operatorname{Lag}_{r}(\mathbb{K}^{2r})\smallsetminus\operatorname{Z}_{P^{-}}$ can be described as follows:
$\mathbb{A}=\left\{\operatorname{Im}\begin{pmatrix}I_{r}\\ X\end{pmatrix}\ \Big{|}\ X\in\operatorname{Mat}_{r}(\mathbb{K}),\ ^{t}\overline{X}=X\right\}.$
Then a computation gives:
$\begin{split}\mathbb{A}\cap\operatorname{Z}_{P}&=\left\{\operatorname{Im}\begin{pmatrix}I_{r}\\ X\end{pmatrix}\ \Big{|}\ ^{t}\overline{X}=X,\ \det(X)=0\right\};\\ \mathbf{C}(P)&=\left\{\operatorname{Im}\begin{pmatrix}I_{r}\\ X\end{pmatrix}\ \Big{|}\ ^{t}\overline{X}=X,\ \det(X)=0\text{ and }^{t}\overline{x}Xx\in\mathbb{R}_{\geq 0}\quad\forall x\in\mathbb{K}^{r}\right\}.\end{split}$
Note that if $r\geq 3$, then the inclusion $\mathbf{C}(P)\subset\mathbb{A}\cap\operatorname{Z}_{P}$ is strict.
3.4.2. The Einstein Universe.
Let $n\geq 2$ and let $(e_{1},\cdots,e_{n+2})$ be the canonical basis of $\mathbb{R}^{n+2}$. For any vector $v\in\mathbb{R}^{n+2}$, we denote by $v_{i}$ the $i$-th coordinate of $v$, that is $v=\sum_{i=1}^{n+2}v_{i}e_{i}$. Let $\mathbf{b}$ be the quadratic form of signature $(n,2)$ on $\mathbb{R}^{n+2}$ defined as:
$\mathbf{b}(v,w)=\sum_{i=1}^{n-1}v_{i}w_{i}-v_{n}w_{n}-\frac{1}{2}(v_{n+1}w_{n+2}+v_{n+2}w_{n+1})\quad\forall v,w\in\mathbb{R}^{n+2}.$
Let $G=\operatorname{SO}(2,n)$ and $P$ be the stabilizer of $[e_{n+2}]$ in $G$. Then $\mathbf{Sb}(G)=G/P$ identifies with the Einstein Universe $\operatorname{Ein}^{n-1,1}$, i.e. the set of isotropic lines of $(\mathbb{R}^{n,2},\mathbf{b})$. Explicitly, the affine chart $\mathbb{A}$ defined by $[e_{n+2}]$ is
$\mathbb{A}=\operatorname{Ein}^{n-1,1}\smallsetminus\mathbb{P}\left(\left(\mathbb{R}e_{n+2}\right)\right)^{\perp_{\mathbf{b}}})=\mathbb{P}\Big{\{}\sum_{i=1}^{n}v_{i}e_{i}+\psi\Big{(}\sum_{i=1}^{n}v_{i}e_{i}\Big{)}e_{n+1}+e_{n+2}\Big{\}},$
where $\psi$ is the quadratic form of signature $(n-1,1)$ on $V:=\operatorname{Span}(e_{1},\cdots,e_{n})$ defined by $\psi\left(\sum_{i=1}^{n}v_{i}e_{i}\right)=\sum_{i=1}^{n-1}v_{i}^{2}-v_{n}^{2}$. The identification $V\simeq\mathbb{A}$ given by
$\sum_{i=1}^{n}v_{i}e_{i}\longmapsto\mathbb{P}\Big{(}\sum_{i=1}^{n}v_{i}e_{i}+\psi\Big{(}\sum_{i=1}^{n}v_{i}e_{i}\Big{)}e_{n+1}+e_{n+2}\Big{)}$
endows $\mathbb{A}$ with the structure of a Minkowski space. Still denoting by $\psi$ the quadratic form induced on $\mathbb{A}$ by this identification, one has:
$\mathbb{A}\cap\operatorname{Z}_{y_{0}}=\{y\in\mathbb{A}\mid\psi(y-y_{0})=0\}=\mathbf{C}(y_{0}),$
see Figure 2. Hence in this case, the inclusion $\mathbf{C}(y_{0})\subset\operatorname{Z}_{y_{0}}\cap\mathbb{A}$ is an equality for all $y_{0}\in\mathbb{A}$. This phenomenon widely simplifies the proof of Theorem 1.4 in the case where $G=\operatorname{SO}(n,2)$; see [CG24].
3.5. Embedding the projective line into $\mathbf{Sb}(G)$
In this section we construct embeddings of the projective line into $\mathbf{Sb}(G)$. The images of these embeddings are what we will define as photons in Section 5. In Lemma 3.8, we investigate the action of $U^{+}$ on these images.
Let $G$ be an $\operatorname{HTT}$ Lie group. We denote by $h_{r}$ the element $h_{\alpha_{r}}$ defined in Section 2.2.3. The root spaces $\mathfrak{g}_{\pm\alpha_{r}}$ are one-dimensional, hence there are unique $v^{\pm}\in\mathfrak{g}_{\pm\alpha_{r}}$ such that
(3.3) $\mathsf{t}_{\mathsf{std}}=(v^{+},h_{r},v^{-})$
is an $\mathfrak{sl}_{2}$-triple, which we call standard. The map $\operatorname{j}_{\mathsf{t}_{\mathsf{std}}}$ associated with $\mathsf{t}_{\mathsf{std}}$ (see Section 2.2.1) induces a group homomorphism $\tau:\operatorname{SL}_{2}(\mathbb{R})\hookrightarrow G$ with kernel contained in {±$\{\pm\operatorname{id}\}$ and with differential $\tau_{*}=\operatorname{j}_{\mathsf{t}_{\mathsf{std}}}$ at $\operatorname{id}$.
Lemma 3.7.
The stabilizer of $P$ in $\operatorname{SL}_{2}(\mathbb{R})$ is the standard Borel subgroup $P_{1}$ of $\operatorname{SL}_{2}(\mathbb{R})$.
  • Proof.
[…]
By Lemma 3.7, the map $\tau$ induces a $\tau$-equivariant embedding $\mathsf{i}:\mathbb{P}(\mathbb{R}^{2})\hookrightarrow\mathbf{Sb}(G)$. It will be convenient to write this map explicitly:
(3.4) $\mathsf{i}([1:t])=\exp(tv^{-})P\quad\forall t\in\mathbb{R}.$
Lemma 3.8.
For any $Y\in\mathfrak{u}^{+}$, there exists $\lambda\in\mathbb{R}$ such that for all $t\in\mathbb{R}\smallsetminus\{-\lambda^{-1}\}$ (with $-\lambda^{-1}=\infty$ if $\lambda=0$), one has:
$\exp(Y)\cdot\mathsf{i}([1:t])=\mathsf{i}\Big{(}\Big{[}1:\frac{t}{1+\lambda t}\Big{]}\Big{)}.$
  • Proof.
    We write $Y=\lambda v^{+}+\sum_{\beta\in\Sigma_{\Theta}^{+}\smallsetminus\{\alpha_{r}\}}Y_{\beta}$ with $Y_{\beta}\in\mathfrak{g}_{\beta}$ for all $\beta\in\Sigma_{\Theta}^{+}\smallsetminus\{\alpha_{r}\}$ and with $\lambda\in\mathbb{R}$. Since $\mathfrak{u}^{+}$ is abelian, one has:
    (3.5) $\exp(Y)\cdot\mathsf{i}([1:t])=\exp(\lambda v^{+})\Bigg{(}\prod_{\beta\in\Sigma_{\alpha_{r}}^{+}\smallsetminus\{\alpha_{r}\}}\exp(Y_{\beta})\Bigg{)}\cdot\mathsf{i}([1:t]).$
    Let $\beta\in\Sigma_{\alpha_{r}}^{+}\smallsetminus\{\alpha_{r}\}$. Since $Y_{\beta}\in\operatorname{Lie}(P)$, one has
    $\exp(Y_{\beta})\cdot\mathsf{i}([1:t])=\exp(Y_{\beta})\exp(tv^{-})P$ $=\exp(Y_{\beta})\exp(tv^{-})\exp(-Y_{\beta})P$
    $=\exp\left(t\operatorname{Ad}\left(\exp\left(Y_{\beta}\right)\right)v^{-}\right)P.$
    By Lemma 3.2, there exists $Z\in\operatorname{Lie}(P)$ such that $[v^{-},Z]=0$. Then
    $\exp\left(t\operatorname{Ad}\left(\exp\left(Y_{\beta}\right)\right)v^{-}\right)P$ $=\exp(t\operatorname{Ad}\left(\exp\left(Y_{\beta}\right)\right)v^{-})P$
    $=\exp(t(v^{-}+Z))P=\exp(tv^{-})\exp(tZ)P=\exp(tv^{-})P,$
    the second last equality holding because $v^{-}$ and $Z$ commute, and the last one holding because $Z\in\operatorname{Lie}(P)$. Hence by induction, Equation (3.5) becomes:
    $\exp(Y)\cdot\mathsf{i}([1:t])$ $=\exp(\lambda v^{+})\cdot\mathsf{i}([1:t])=\tau(e^{\lambda\operatorname{E}})\cdot\mathsf{i}([1:t])$
    $=\mathsf{i}(e^{\lambda\operatorname{E}}\cdot[1:t])=\mathsf{i}\Big{(}\Big{[}1:\frac{t}{1+\lambda t}\Big{]}\Big{)},$
[…]
3.6. Proximal representations of $\operatorname{HTT}$ Lie groups
We denote by $\omega_{r}$ the fundamental weight $\omega_{\alpha_{r}}$ associated with the root $\alpha_{r}$, with the notation of Section 2.2.3. There always exists $N\in\mathbb{N}^{*}$ such that $G$ admits a finite-dimensional real irreducible linear representation of highest weight $N\omega_{r}$ (see e.g. [GGKW17, Lem. 3.2]). Such a representation is automatically $\{\alpha_{r}\}$-proximal. We will use the following classical lemma throughout this paper:
Lemma 3.9.
Let $G$ be an $\operatorname{HTT}$ Lie group and let $(V,\rho)$ be a finite-dimensional real irreducible linear representation of $G$ with highest weight $\chi=N\omega_{r}$ for some $N\in\mathbb{N}_{>0}$. Let $e_{1}\in V^{\chi}\smallsetminus\{0\}$. Then $\rho_{*}(v^{-})^{k}e_{1}\neq 0$ for all $k\leq N$, and $\rho_{*}(v^{-})^{k}e_{1}=0$ for all $k\geq N+1$.
  • Proof.
    By the definition of $\omega_{r}$ one has $0\neq\rho_{*}(v^{-})^{k}e_{1}\in V^{N\omega_{r}-k\alpha_{r}}$ for all $0\leq k\leq N$ (see e.g. [GW09, Lem. 3.2.9]).
    Since $\operatorname{\Delta}$ is of type $C_{r}$, the Weyl group $W$ acts by signed permutations of the $(\varepsilon_{i})_{1\leq i\leq r}$ (see e.g. [Kna96]) and $\omega_{r}=\sum_{i=1}^{r}\varepsilon_{i}$. Hence the convex envelope of the orbit $W\cdot(N\omega_{r})$ of $N\omega_{r}$ by $W$ in $\mathfrak{a}$ is $C:=\{N\sum_{i=1}^{r}\delta_{i}\varepsilon_{i}\ |\ \delta_{i}\in\{\pm 1\}\}$.
    The set of weights of $(V,\rho)$ is the intersection of $C$ with the translate $(N\omega_{r})+\sum_{i=1}^{r}\mathbb{Z}\alpha_{i}$ of the root lattice (see e.g. [GW09, Prop. 3.2.10]). For $k\geq N+1$, one has
    $N\omega_{r}-k\alpha_{r}=N\sum_{i=1}^{r-1}\varepsilon_{i}-(2k-N)\varepsilon_{r}\notin C$
[…]
4. Proper domains and their automorphism group
In this section we recall some definitions and properties of domains in a flag manifold. In particular, we introduce the diamonds (see Definition 4.6) in $\mathbf{Sb}(G)$, and recall their basic properties. These proper domains are quasi-homogeneous, and Theorem 1.4 states that they are the only proper domains in $\mathbf{Sb}(G)$ with this property.
4.1. Generalities on proper domains
The notions recalled in this section are generalizations of classical notions of convex projective geometry, and most of them were introduced in [Zim18a].
Let $G$ be a noncompact simple Lie group and $\Theta\subset\operatorname{\Delta}$ a subset of the simple roots.
Definition 4.1.
Let $\Omega\subset\mathscr{F}_{\Theta}$ be an open subset. We say that $\Omega$ is:
  1. (1)
    a domain if $\Omega$ is nonempty and connected;
  2. (2)
    proper if there exists $\xi\in G/P^{-}_{\Theta}$ such that $\overline{\Omega}\cap\operatorname{Z}_{\xi}=\emptyset$. In particular, if $\overline{\Omega}\cap\operatorname{Z}_{P_{\Theta}^{-}}=\emptyset$ then we will say that $\Omega$ is proper in $\mathbb{A}$. This is equivalent to saying that $\overline{\Omega}\subset\mathbb{A}$.
Remark 4.2.
Given a proper domain $\Omega$ of $\mathscr{F}_{\Theta}$, we will always be able to assume that $\Omega$ is proper in $\mathbb{A}$. Indeed, since $G$ acts transitively on $\mathscr{F}_{\Theta}^{-}$, there exists $g\in G$ such that $\overline{g\cdot\Omega}\subset\mathbb{A}$, and the properties we will investigate on $\Omega$ will be invariant under the action of $G$ on $\mathscr{F}_{\Theta}$. In this case, it will be possible to see $\Omega$ as a bounded domain of the affine space $\mathbb{A}$.
4.1.1. The automorphism group.
Given an open subset $\Omega\subset\mathscr{F}_{\Theta}$, we define its automorphism group
$\mathsf{Aut}(\Omega)=\left\{g\in G\mid g\cdot\Omega=\Omega\right\}.$
By [Zim18a], the group $\mathsf{Aut}(\Omega)$ is a Lie subgroup of $G$. Moreover, it acts properly discontinuously on $\Omega$ as soon as $\Omega$ is a proper domain.
Remark 4.3.
In the case where $G$ is an $\operatorname{HTT}$ Lie group and $\mathscr{F}_{\Theta}=\mathbf{Sb}(G)$, the group $\mathsf{Aut}(\Omega)$ is commensurable to the conformal group of $\Omega$, that is, the group of all the invertible maps from $\Omega$ to itself that preserve the causal structure of $\Omega$ [Kan11, Thm 2.3].
A domain $\Omega$ is said to be quasi-homogeneous if there exists a compact subsets $\mathscr{K}\subset\Omega$ such that $\Omega=\mathsf{Aut}(\Omega)\cdot\mathscr{K}$. Note that if $G$ is an $\operatorname{HTT}$ Lie group, then this notion only depends on the Lie algebra $\mathfrak{g}$ of $G$, because $\operatorname{Ad}(G)$ has finite index in $\mathsf{Aut}(\mathfrak{g})$ (see e.g. [Sat14]). The domain $\Omega$ is said to be symmetric if for any $x\in\Omega$ there exists an order-two element $s_{x}\in\mathsf{Aut}_{1}(\mathfrak{g})$ such that $s_{x}\cdot\Omega=\Omega$ and $x$ is the only fixed point of $s_{x}$ in $\Omega$. This notion only depends on the Lie algebra $\mathfrak{g}$ of $G$ (recall that, if $G$ is an $\operatorname{HTT}$ Lie group, then one has $\mathsf{Aut}_{1}(\mathfrak{g})=\mathsf{Aut}(\mathfrak{g})$ — see Section 3.1).
Note that $\mathsf{Aut}(g\cdot\Omega)=g\mathsf{Aut}(\Omega)g^{-1}$ for all $g\in G$; therefore the property of being quasi-homogeneous, resp. symmetric, is invariant under the action of $G$ on $\mathscr{F}_{\Theta}$. In our case, i.e. when $G$ is an $\operatorname{HTT}$ Lie group, by Remark 4.2 and Section 3.2, it will always be possible to assume that $\Omega$ is proper in $\mathbb{A}$, and given a point $x\in\overline{\Omega}$, we can always assume that $x=P$, up to translating $\Omega$ by an element of $G$.
We will say that a metric $d$ on $\Omega$ is $\mathsf{Aut}(\Omega)$-invariant if $d(g\cdot x,g\cdot y)=d(x,y)$ for all $x,y\in\Omega$ and $g\in\mathsf{Aut}(\Omega)$.
4.1.2. The dual.
Let $\Omega\subset\mathscr{F}_{\Theta}$ be a domain. The dual of $\Omega$ is the set
$\Omega^{*}:=\{\xi\in G/P_{\Theta}^{-}\mid\operatorname{Z}_{\xi}\cap\Omega=\emptyset\}\subset\mathscr{F}_{\Theta}^{-}.$
Let us recall some properties of this set (see [Zim18a]):
  1. (1)
    The set $\Omega^{*}$ is $\mathsf{Aut}(\Omega)$-invariant, i.e. one has $\mathsf{Aut}(\Omega)\subset\mathsf{Aut}(\Omega^{*})$.
  2. (2)
    If $\Omega$ is open, then $\Omega^{*}$ is compact.
  3. (3)
    The domain $\Omega$ is proper if, and only if, its dual $\Omega^{*}$ has nonempty interior.
The domain $\Omega$ is dually convex, if for any $p\in\partial\Omega$ there exists $\xi\in\Omega^{*}$ such that $p\in\operatorname{Z}_{\xi}$. One has:
Proposition 4.4.
([Zim18a, Cor. 9.3]) Any proper quasi-homogeneous domain of $\mathscr{F}_{\Theta}$ is dually convex.
Remark 4.5.
Dual convexity is a generalization to arbitrary flag manifolds of the classical notion of convexity in the projective space, using the dual characterization of convexity. Proposition 4.4 is then analogous to a classical result of Kobayashi [Sho84].
4.2. Diamonds
In this section, let $G$ be an $\operatorname{HTT}$ Lie group. Given two transverse points $p,q\in\mathbf{Sb}(G)$, the set $\mathbf{Sb}(G)\smallsetminus(\operatorname{Z}_{p}\cup\operatorname{Z}_{q})$ admits several connected components, exactly two of which are proper.
Definition 4.6.
A subset $\Omega$ of $\mathbf{Sb}(G)$ is called a diamond if there exist a (unique) pair of transverse points $p,q\in\mathbf{Sb}(G)$ such that $\Omega$ is one of the two proper connected components of $\mathbf{Sb}(G)\smallsetminus(\operatorname{Z}_{p}\cup\operatorname{Z}_{q})$. The two points $p,q$ are then called the endpoints of $\Omega$.
Let $\mathbf{D}_{\mathsf{std}}:=\mathbf{I}^{+}(P)$. Recall the order-two element $k_{0}\in G$ such that $k_{0}Pk_{0}^{-1}=P^{-}$ (see Section 2.3). Then $\mathbf{D}_{\mathsf{std}}^{\prime}:=\mathbf{I}^{-}(P)=k_{0}\cdot\mathbf{D}_{\mathsf{std}}$ is the interior of the dual of $\mathbf{D}_{\mathsf{std}}$ (see e.g. [GW22, Lem. 13.11]), and the domains $\mathbf{D}_{\mathsf{std}}$ and $\mathbf{D}_{\mathsf{std}}^{\prime}$ are exactly the two diamonds with endpoints $P$ and $P^{-}$. They are proper in $\mathbf{Sb}(G)$ — although they are not proper in $\mathbb{A}$.
Given two transverse points $p,q\in\mathbf{Sb}(G)$, one has $(p,q)=g\cdot(P,P^{-})$ for some $g\in G$. The two diamonds $g\cdot\mathbf{D}_{\mathsf{std}}$ and $g\cdot\mathbf{D}_{\mathsf{std}}^{\prime}$ are the diamonds with endpoints $p$ and $q$.
By the two previous paragraphs, any diamond is a $G$-translate of $\mathbf{D}_{\mathsf{std}}$. In particular, up to the action of $G$ on $\mathbf{Sb}(G)$, there is only one model of diamond in $\mathbf{Sb}(G)$. It is convenient to consider models of diamonds that are proper in $\mathbb{A}$:
Definition 4.7.
If $p,q\in\mathbb{A}$ and $q\in\mathbf{I}^{+}(p)$, we define $\mathbf{D}(p,q)$ as the set $\mathbf{I}^{+}(p)\cap\mathbf{I}^{-}(q)$. It is one of the two diamonds with endpoints $p$ and $q$.
For $p,q\in\mathbb{A}$ and $q\in\mathbf{I}^{+}(p)$, the diamond $\mathbf{D}(p,q)$ is the only one of the two diamonds with endpoints $p$ and $q$ that is proper in $\mathbb{A}$; see Figure 2.
The following fact is well-known (see e.g. [GW16, Prop. 3.7, 5.2 and Remark 5.4] and [Kan11, Thm 2.3 and 3.5]):
Fact 4.8.
The automorphism group of $\mathbf{D}_{\mathsf{std}}$ in $\mathsf{Aut}(\mathfrak{g})$ contains a symmetry. The diamond $\mathbf{D}_{\mathsf{std}}$ is thus a symmetric domain of $\mathbf{Sb}(G)$.
Moreover, the action of the identity component $L^{0}$ of $L$ on $\mathbf{D}_{\mathsf{std}}$ is transitive and the stabilizer of a point in $\mathbf{D}_{\mathsf{std}}$ is a maximal compact subgroup of $L^{0}$, so that any diamond is a model for the symmetric space of $L^{0}$.
Since $L$ has one-dimensional center, there exist finite-index subgroups $H_{1}\leq\mathsf{Aut}(\mathbf{D}_{\mathsf{std}})$ and $H_{2}\leq\mathrm{Isom}(\mathbb{R}\times X_{L_{s}})$ such that the diamond $\mathbf{D}_{\mathsf{std}}$ is $(H_{1},H_{2})$-equivariantly diffeomorphic to ×$\mathbb{R}\times X_{L_{s}}$, where $X_{L_{s}}$ is the symmetric space of $L_{s}$ (see e.g. [Kan06, Table I]). These identifications are listed in Table 2.
Remark 4.9.
  1. (1)
    The family of diamonds $\mathbf{D}(p,q)$ with $p,q\in\mathbb{A}$ and $q\in\mathbf{I}^{+}(p)$ forms a basis of neighborhoods of $\mathbb{A}$.
  2. (2)
    Since $L_{s}$ admits cocompact lattices [Bor63], any diamond is divisible, in the sense that there exists a discrete subgroup $\Gamma\leq\mathsf{Aut}(\Omega)$ and a compact subset $\mathscr{K}\subset\Omega$ such that $\Omega=\Gamma\cdot\mathscr{K}$.
  3. (3)
    Diamonds are defined more generally in any flag manifold $\mathscr{F}_{\Theta}$ admitting a $\Theta$-positive structure [GLW21]. By Zimmer’s theorem (Fact 1.3) and Fact 4.8, these diamonds are quasi-homogeneous (resp. divisible) if, and only if, the flag manifold $\mathscr{F}_{\Theta}$ is the Shilov boundary associated with an $\operatorname{HTT}$ Lie group.
Example 4.10 (Explicit construction of diamonds).
Let us see what diamonds look like, for different values of $G$.
(1) $G=\operatorname{SO}(n,2)$ with $n\geq 3$. Given two transverse points $p,q\in\operatorname{Ein}^{n-1,1}$, the set $\mathbf{Sb}(G)\smallsetminus(\operatorname{Z}_{p}\cup\operatorname{Z}_{q})$ has exactly three connected components (see Figure 2 for $p=P$ and $q=P^{-}$). For a general $\operatorname{HTT}$ Lie group $G$, the set $\mathbf{Sb}(G)\smallsetminus(\operatorname{Z}_{p}\cup\operatorname{Z}_{q})$, where $p,q$ are two transverse points, may have more connected components. If $\dim(\mathbf{Sb}(G))\geq 3$, then there are exactly $(r+1)$ connected components, where $r$ is the real rank of $G$ (see e.g. [Kan98]). This is related to the fact that, in general, the inclusion $\mathbf{C}(P)\subset\operatorname{Z}_{P}\cap\mathbb{A}$ is strict (see Section 3.4.1), and is an obstruction to a direct generalization to any $\operatorname{HTT}$ Lie group of the proof of Theorem 1.4 for $G=\operatorname{SO}(n,2)$ ($n\geq 2$) established in [CG24].
(2) $G=\operatorname{Sp}(2r,\mathbb{R})$. Let us take the notation of Section 3.4. Then
$\mathbf{D}_{\mathsf{std}}=\Big{\{}\operatorname{Im}\begin{pmatrix}I_{r}\\ X\end{pmatrix}\mid X\in\operatorname{Sym}_{r}^{++}(\mathbb{R})\Big{\}};\ \mathbf{D}_{\mathsf{std}}^{\prime}=\Big{\{}\operatorname{Im}\begin{pmatrix}I_{r}\\ -X\end{pmatrix}\mid X\in\operatorname{Sym}_{r}^{++}(\mathbb{R})\Big{\}},$
where $\operatorname{Sym}_{r}^{++}(\mathbb{R})$ is the set of positive definite symmetric matrices of size $r\leavevmode\nobreak\ \times\leavevmode\nobreak\ r$. In particular, the diamond $\mathbf{D}_{\mathsf{std}}$ is a model for the symmetric space ×$\mathbb{R}\times\left(\operatorname{SL}(r,\mathbb{R})/\operatorname{SO}(r)\right)$. The analogue diamonds for other Shilov boundaries associated with $\operatorname{HTT}$ Lie groups are given in Table 2 (see also Table 1 of [GW16]).
$P$$\mathbf{I}^{+}(P)=\mathbf{D}_{\mathsf{std}}$$\mathbf{I}^{-}(P)=\mathbf{D}_{\mathsf{std}}^{\prime}$
Figure 1. Past and future of $P$ in $\mathbb{A}$ in the case where $\mathfrak{g}=\mathfrak{so}(3,2)$
$q$$p$
Figure 2. The diamond $\mathbf{D}(p,q)$ for $p,q\in\mathbb{A}$ and $q\in\mathbf{I}^{+}(p)$ (greyed-out area), seen in $\mathbb{A}\simeq\mathbb{R}^{2,1}$ for $\mathfrak{g}=\mathfrak{so}(3,2)$.
$G$ $\mathbf{D}_{\mathsf{std}}\simeq$
$\operatorname{SO}(n,2)$$\mathbb{R}\times\mathbb{H}^{n}$
$\operatorname{Sp}(2r,\mathbb{R})$$\mathbb{R}\times(\operatorname{SL}(r,\mathbb{R})/\operatorname{SO}(r))$
$\operatorname{SU}(r,r)$$\mathbb{R}\times(\operatorname{SL}(r,\mathbb{C})/\operatorname{SU}(r))$
$\operatorname{SO}^{*}(4r)$$\mathbb{R}\times(\operatorname{SL}(r,\mathbb{H})/\operatorname{Sp}(r))$
$E_{7(-25)}$$\mathbb{R}\times(E_{6(-26)}/F_{4})$
Table 2. The diamonds in $\mathbf{Sb}(G)$ for every $\operatorname{HTT}$ Lie group (up to local isomorphism).
5. Photons
In this section, we introduce the notion of photon in the Shilov boundary $\mathbf{Sb}(G)$ associated with an $\operatorname{HTT}$ Lie group $G$. The name refers to the well-known photons of the Einstein Universe $\operatorname{Ein}^{n-1,1}$. We will use these generalized photons to define the Kobayashi pseudo-metric of a domain of $\mathbf{Sb}(G)$ in Section 6.
5.1. Definition and basic properties
Let $G$ be an $\operatorname{HTT}$ Lie group. Recall the element $h_{r}\in\mathfrak{a}$ defined in Section 3.5. An $\mathfrak{sl}_{2}$-triple $\mathsf{t}=(e,h,f)$ in $\mathfrak{g}$ is said to be photon-generating if $h$ is conjugate to $h_{r}$ in $G$, i.e. if there exists $g\in G$ such that $h=\operatorname{Ad}(g)h_{r}$. The embedding $\operatorname{j}_{\mathsf{t}}$ defined in Section 2.2.1 induces an action of $\operatorname{SL}_{2}(\mathbb{R})$ on $\mathbf{Sb}(G)$, referred to as a photon-generating action. This action has a unique closed orbit, which is a topological circle (whose elements are all the attracting fixed points of proximal elements of $\operatorname{SL}_{2}(\mathbb{R})$ for this action).
Definition 5.1.
A photon is the unique closed orbit of a photon-generating action of $\operatorname{SL}_{2}(\mathbb{R})$ on $\mathbf{Sb}(G)$. We denote by $\mathscr{L}$ the set of all photons of $\mathbf{Sb}(G)$.
We denote by $\Lambda_{\mathsf{std}}$ the standard photon defined by the standard $\mathfrak{sl}_{2}$-triple $\mathsf{t}_{\mathsf{std}}$ of (3.3). The map $\mathsf{i}$ of (3.4) is then a parametrization of $\Lambda_{\mathsf{std}}$. We will mostly use the following equality in the rest of this paper:
=$\mathscr{L}=\{g\cdot\Lambda_{\mathsf{std}}\mid g\in G\}.$
The following lemma follows from Lemma 3.8:
Lemma 5.2.
One has $U^{+}\cdot\Lambda_{\mathsf{std}}=\Lambda_{\mathsf{std}}$.
Remark 5.3.
Recall that a Schubert subvariety of $\mathbf{Sb}(G)$ is the Zariski closure in $\mathbf{Sb}(G)$ of a subset of the form $X(w)=P_{\operatorname{\Delta}}w\cdot P$, where $\omega$ is an element of the Weyl group of $\Sigma$ stabilizing $\alpha_{r}$ (see e.g. [Kna96]). Noticing that $\Lambda_{\mathsf{std}}$ is a one-dimensional Schubert subvariety of $\mathbf{Sb}(G)$ gives a direct proof of Lemma 5.2.
5.1.1. Parametrization of a photon
Given a photon $\Lambda\in\mathscr{L}$, there exists $g\in G$ such that $\Lambda=g\cdot\Lambda_{\mathsf{std}}$. The map $\mathsf{i}_{g}$ defined as
(5.1) $\mathsf{i}_{g}:\begin{cases}\mathbb{P}(\mathbb{R}^{2})&\longrightarrow\mathbf{Sb}(G)\\ x&\longmapsto g\cdot\mathsf{i}(x)\end{cases}$
is then a parametrization of $\Lambda$. A priori, this parametrization depends on the choice of $g\in G$ such that $\Lambda=g\cdot\Lambda_{\mathsf{std}}$ (although its image does not). The next lemma shows that two parametrizations given by different choices of $g\in G$ such that $\Lambda=\operatorname{Ad}(g)\cdot\Lambda_{\mathsf{std}}$ only differ by a projective reparametrization of $\mathbb{P}(\mathbb{R}^{2})$.
Lemma 5.4.
One has
$\operatorname{Stab}_{G}(\Lambda_{\mathsf{std}})=\tau(\operatorname{SL}_{2}(\mathbb{R}))\times\operatorname{Cent}_{G}\left(\tau\left(\operatorname{SL}_{2}(\mathbb{R})\right)\right),$
where $\operatorname{Stab}_{G}(\Lambda_{\mathsf{std}})$ is the stabilizer of $\Lambda_{\mathsf{std}}$ in $G$ and $\operatorname{Cent}_{G}\left(\tau\left(\operatorname{SL}_{2}(\mathbb{R})\right)\right)$ is the centralizer in $G$ of the group $\tau(\operatorname{SL}_{2}(\mathbb{R}))$.
By the definition of $\Lambda_{\mathsf{std}}$, the centralizer of $\tau(\operatorname{SL}_{2}(\mathbb{R}))$ acts trivially on $\Lambda_{\mathsf{std}}$. Therefore Lemma 5.4 implies that $\operatorname{Stab}_{G}(\Lambda_{\mathsf{std}})$ acts on $\Lambda_{\mathsf{std}}$ by projective transformations. For any photon $\Lambda$, the set $\Lambda$ is a $G$-translate of $\Lambda_{\mathsf{std}}$, so $\operatorname{Stab}_{G}(\Lambda)$ acts on $\Lambda$ by projective transformations.
  • Proof of Lemma 5.4.
    The inclusion $\tau(\operatorname{SL}_{2}(\mathbb{R}))\times\operatorname{Cent}_{G}\left(\tau\left(\operatorname{SL}_{2}(\mathbb{R})\right)\right)\subset\operatorname{Stab}_{G}(\Lambda_{\mathsf{std}})$ follows from the definition of $\Lambda_{\mathsf{std}}$. Let us prove the converse inclusion.
Let $g\in\operatorname{Stab}_{G}(\Lambda_{\mathsf{std}})$. First assume that $gP\in\mathbb{A}$. Then, by (2.4), we can write $g=\exp(X)\ell\exp(Y)$ with $Y\in\mathfrak{u}^{+}$, $X\in\mathfrak{u}^{-}$ and $\ell\in L$. Since $g$ stabilizes $\Lambda_{\mathsf{std}}$, there exist $\mu,\delta\in\mathbb{R}$ such that $\ell=\exp(\mu h_{r})$ and $X=\delta v^{-}$. By Lemma 3.8, there exists $\lambda\in\mathbb{R}$ such that for all $t\in\mathbb{R}\smallsetminus\left\{-\lambda^{-1}\right\}$, one has
$g\cdot\mathsf{i}([1:t])=\exp(X)\exp(\mu h_{r})\cdot\mathsf{i}\Big{(}\Big{[}1:\frac{t}{1+\lambda t}\Big{]}\Big{)}=\exp(X)\exp(\mu h_{r})\cdot\mathsf{i}\left(e^{\lambda\operatorname{F}}\left[1:z\right]\right).$
Moreover, one has $\exp(X)=\tau\left(e^{\delta\operatorname{E}}\right)$ and $\ell=\tau\left(e^{\mu\operatorname{H}}\right)$. Then, by $\tau$-equivariance of $\mathsf{i}$, one has $g\cdot\mathsf{i}([1:t])=\mathsf{i}\left(A\cdot\left[1:t\right]\right)$ with $A=e^{\delta\operatorname{F}}e^{\mu\operatorname{H}}e^{\lambda\operatorname{E}}\in\operatorname{SL}_{2}(\mathbb{R})$. By continuity, this equality holds for all $x\in\mathbb{P}(\mathbb{R}^{2})$. In particular, the element $\tau(A)^{-1}g$ fixes every point of $\Lambda_{\mathsf{std}}$. By the definition of $\Lambda_{\mathsf{std}}$, this implies that $\tau(A)^{-1}g\in\operatorname{Cent}_{G}\left(\tau\left(\operatorname{SL}_{2}(\mathbb{R})\right)\right)$.
Now if $gP\notin\mathbb{A}$, since $g$ preserves $\Lambda_{\mathsf{std}}$ one must have $gP=\mathsf{i}([0:1])$. Since $\operatorname{SL}_{2}(\mathbb{R})$ acts transitively on $\mathbb{P}(\mathbb{R}^{2})$, there exists $B\in\operatorname{SL}_{2}(\mathbb{R})$ such that
$gP=\mathsf{i}([0:1])=\mathsf{i}(B\cdot[1:0])=\tau(B)\mathsf{i}([1:0]),$
so that $\tau(B)^{-1}g\in\operatorname{Stab}_{G}(\Lambda_{\mathsf{std}})$ satisfies $\tau(B)^{-1}gP=P\in\mathbb{A}$. Then by the previous case, one has $\tau(B)^{-1}g\in\tau(\operatorname{SL}_{2}(\mathbb{R}))\times\operatorname{Cent}_{G}\left(\tau\left(\operatorname{SL}_{2}(\mathbb{R})\right)\right)$. Hence
$g=\tau(B)\tau(B)^{-1}g\in\tau(\operatorname{SL}_{2}(\mathbb{R}))\times\operatorname{Cent}_{G}\left(\tau\left(\operatorname{SL}_{2}(\mathbb{R})\right)\right)\quad\hfill\square$
5.1.2. Photons in the standard affine chart.
The following lemma states that photons intersecting $\mathbb{A}$ are compactifications of certain affine lines of $\mathbb{A}$.
Lemma 5.5.
Let $\Lambda\in\mathscr{L}$. If $\Lambda\cap\mathbb{A}$ is nonempty, then it is an affine line in $\mathbb{A}$, and $\Lambda\cap\operatorname{Z}_{P^{-}}$ is a singleton.
  • Proof.
    Assume that $\Lambda\cap\mathbb{A}\neq\emptyset$. Let $g\in G$ be such that $\Lambda=g\cdot\Lambda_{\mathsf{std}}$. Since $\Lambda$ is not contained in $\operatorname{Z}_{P^{-}}$, there exists $t\in\mathbb{R}$ such that $g\exp(tv^{-})P\in\mathbb{A}$. Since $\exp(tv^{-})$ stabilizes $\Lambda_{\mathsf{std}}$, we have $g\exp(tv^{-})\cdot\Lambda_{\mathsf{std}}=\Lambda$. Hence, to replacing $g$ with $g\exp(tv^{-})$, we may assume that $gP\in\mathbb{A}$.
    By (2.4), one can thus write $g=\exp(X)\ell\exp(Y)$ with $X\in\mathfrak{u}^{-}$, $Y\in\mathfrak{u}^{+}$ and $\ell\in L$. By Lemma 5.2, one then has $g\cdot\Lambda_{\mathsf{std}}=\exp(X)\ell\cdot\Lambda_{\mathsf{std}}$. Hence
    (5.2) $\varphi_{\mathsf{std}}^{-1}\left(\Lambda\cap\mathbb{A}\right)=\varphi_{\mathsf{std}}^{-1}\left((g\cdot\Lambda_{\mathsf{std}})\cap\mathbb{A}\right)=X+\operatorname{Ad}(\ell)\mathfrak{g}_{-\alpha_{r}}\subset\mathfrak{u}^{-}$
    is an affine line of $\mathfrak{u}^{-}$. Hence $\Lambda\cap\mathbb{A}$ is an affine line of $\mathbb{A}$ for the canonical affine structure.
[…]
Lemma 5.5 implies that for all $\xi\in\mathbf{Sb}(G)$ such that $\Lambda_{\mathsf{std}}\not\subset\operatorname{Z}_{\xi}$, the set $\Lambda_{\mathsf{std}}\cap\operatorname{Z}_{\xi}$ is a singleton.
Remark 5.6.
(1) It follows directly from the definition of $c^{0}$ that $(\Lambda\cap\mathbb{A})\subset\mathbf{C}(x)$ for any $x\in\mathbb{A}$ and $\Lambda\in\mathscr{L}$ such that $x\in\Lambda$. Moreover, given a point $x\in\mathbf{Sb}(G)$, we have the equivalence: $\Lambda_{\mathsf{std}}\subset\operatorname{Z}_{x}$ if, and only if, $x\in\Lambda_{\mathsf{std}}$.
(2) More generally, given any $\mathfrak{sl}_{2}$-triple of $\mathfrak{g}$, the unique closed orbit of the induced action of $\operatorname{SL}_{2}(\mathbb{R})$ on $\mathbf{Sb}(G)$ is a topological circle, sometimes simply called a circle. In [GLW21], positive circles are defined from triples $\mathsf{t}=(e,h,f)$ where $f$ is conjugate to an element of $c^{0}$. Here we state their definition only in the context of Shilov boundaries of $\operatorname{HTT}$s, but positive circles are defined in any flag manifold admitting a $\Theta$-positive structure and are particular magical $\mathfrak{sl}_{2}$-triples [BCGP+].
(3) The definition of a positive circle [GLW21] is in some sense opposite to that of a photon, and its properties are opposite to those of photons (for instance, the centralizer of the associate embedding of $\operatorname{SL}_{2}(\mathbb{R})$ into $G$ is compact).
(4) Photons have already been defined in [LZ19] for the Grassmannians $\operatorname{Gr}_{p}(\mathbb{R}^{n})$, where they are called rank-one lines. For the natural identification of $\operatorname{Lag}_{r}(\mathbb{R}^{2r})$ with a submanifold of $\operatorname{Gr}_{r}(\mathbb{R}^{2r})$, the photons of $\operatorname{Lag}_{r}(\mathbb{R}^{2r})$ are exactly the rank-one lines of $\operatorname{Gr}_{r}(\mathbb{R}^{2r})$ that are contained in $\operatorname{Lag}_{r}(\mathbb{R}^{2r})$.
5.2. Intersection polynomials
In this section we define the intersection polynomials, which algebraically describe the intersection between the standard photon $\Lambda_{\mathsf{std}}$ and the nontransverse set $\operatorname{Z}_{\xi}$ of a point $\xi\in\mathbf{Sb}(G)$ such that $\Lambda_{\mathsf{std}}\not\subset\operatorname{Z}_{\xi}$.
Recall the notation of Sections 2.4 and 3.6. Fix a finite-dimensional real irreducible linear representation $(V,\rho)$ of $G$ with highest weight $\chi=N\omega_{r}$ for some $N\in\mathbb{N}_{>0}$. We let $\iota:\mathbf{Sb}(G)\hookrightarrow\mathbb{P}(V)$ and $\iota^{*}:\mathbf{Sb}(G)\hookrightarrow\mathbb{P}(V^{*})$ be the two embeddings induced by $\rho$ by Proposition 2.1, and we fix a vector $e_{1}\in V^{\chi}\smallsetminus\{0\}$. By Lemma 3.9 and Equation (3.4), we have, for $t\in\mathbb{R}$,
(5.3) $\iota\circ\mathsf{i}([1:t])=[e^{t\rho_{*}(v^{-})}\cdot e_{1}]=\big{[}e_{1}+t\rho_{*}(v^{-})e_{1}+\cdots+\frac{t^{N}}{N!}\rho_{*}(v^{-})^{N}e_{1}\big{]}$
Let us define the dense open subset
$\mathscr{X}:=\mathbf{Sb}(G)\smallsetminus\Lambda_{\mathsf{std}}$
of $\mathbf{Sb}(G)$. By the “moreover” part of Remark 5.6.(1), we have $\mathscr{X}=\left\{\xi\in\mathbf{Sb}(G)\mid\Lambda_{\mathsf{std}}\not\subset\operatorname{Z}_{\xi}\right\}$. Given $\xi\in\mathscr{X}$, we choose any lift $f\in V^{*}\smallsetminus\{0\}$ of $\iota^{*}(\xi)\in\mathbb{P}(V^{*})$. Since $\Lambda_{\mathsf{std}}\not\subset\operatorname{Z}_{\xi}$, by Proposition 2.1, the polynomial defined by
$f\big{(}e^{t\rho_{*}(v^{-})}\cdot e_{1}\big{)}=f(e_{1})+tf(\rho_{*}(v^{-})e_{1})+\cdots+t^{N}f(\rho_{*}(v^{-})^{N}e_{1})\quad\forall t\in\mathbb{R}$
is nonzero. Then there exists a maximal $0\leq n(\xi)\leq N$ such that $f(\rho_{*}(v^{-})^{n(\xi)}e_{1})\neq 0$, and $n(\xi)$ does not depend on the choice of the lift $f$ of $\iota^{*}(\xi)$. Hence we may choose $f$ such that $f(\rho_{*}(v^{-})^{n(\xi)}e_{1})=n(\xi)!$. This defines a polynomial $\mathsf{Q}_{\xi}^{\rho}$ with coefficients in $\mathbb{R}$, depending only on $\xi$ and $(V,\rho)$:
(5.4) $\mathsf{Q}_{\xi}^{\rho}(t)=f\big{(}e^{t\rho_{*}(v^{-})}\cdot e_{1}\big{)}=f(e_{1})+tf(\rho_{*}(v^{-})e_{1})+\cdots+t^{n(\xi)}\quad\forall t\in\mathbb{R}.$
Definition 5.7.
Given a point $\xi\in\mathscr{X}$, the polynomial $\mathsf{Q}_{\xi}^{\rho}$ defined in (5.4) is called the intersection polynomial of $\xi$ associated with the representation $(V,\rho)$.
Now let
(5.5) $\mathscr{A}_{\rho}:=\left\{\xi\in\mathbf{Sb}(G)\mid[\rho_{*}(v^{-})^{N}e_{1}]\notin\iota^{*}(\xi)\right\}\subset\mathscr{X}$
be the set of all elements of $\mathscr{X}$ such that $n(\xi)$ is maximal.
Lemma 5.8.
The set $\mathscr{A}_{\rho}$ is open and dense in $\mathbf{Sb}(G)$.
  • Proof.
    By irreducibility of $\rho$, for any open set $\mathscr{O}\subset\mathbf{Sb}(G)$, there exist $x_{1},...,x_{D}\in\mathscr{O}$ such that $\iota^{*}(\xi_{1})\oplus...\oplus\iota^{*}(\xi_{D})=V^{*}$ (see e.g. [Zim18a, Lem. 4.7]). If the set
    $F:=\{\xi\in\mathbf{Sb}(G)\mid[\rho_{*}(v^{-})^{N}e_{1}]\in\iota^{*}(\xi)\}$
[…]
Let $\xi\in\mathscr{X}$ and $t\in\mathbb{R}$. One has
$\mathsf{Q}_{\xi}^{\rho}(t)=0$ $\Longleftrightarrow\ f\big{(}e^{t\rho_{*}(v^{-})}\cdot e_{1}\big{)}\Longleftrightarrow\ \iota(\exp(tv^{-})P)\in\iota^{*}(\xi)$
$\Longleftrightarrow\ \exp(tv^{-})P\in\operatorname{Z}_{\xi}\Longleftrightarrow\ \mathsf{i}([1:t])=\exp(tv^{-})P\in\Lambda_{\mathsf{std}}\cap\operatorname{Z}_{\xi},$
the last equivalence holding by (3.4) and Proposition 2.1. Hence the real roots of $\mathsf{Q}_{\xi}^{\rho}$ describe the intersection points of $\Lambda_{\mathsf{std}}\cap\mathbb{A}$ with $\operatorname{Z}_{\xi}$. By Lemma 5.5, the set $\Lambda_{\mathsf{std}}\cap\operatorname{Z}_{\xi}$ is a singleton, so the polynomial $\mathsf{Q}_{\xi}^{\rho}$ has at most one real root $t$, satisfying $\mathsf{i}([1:t])\in\operatorname{Z}_{\xi}$. We will see in Section 6.2.2 that the complex roots of $\mathsf{Q}_{\xi}^{\rho}$ also describe the intersection of two sets, corresponding to complexifications of $\Lambda_{\mathsf{std}}\cap\mathbb{A}$ and $\operatorname{Z}_{\xi}$. This is why we call the polynomial $\mathsf{Q}_{\xi}^{\rho}$ the intersection polynomial of $\xi$.
6. An invariant metric
In this section we define the Kobayashi metric of a proper domain $\Omega$ of $\mathbf{Sb}(G)$. Constructions of Kobayashi metrics are classical and were initiated by S. Kobayashi [Kob67, Sho84]. The properties of the Kobayashi metric (in particular, its properness, see Corollary 6.13) in the quasi-homogeneous case will allow us to relate the geometry of the boundary of a proper quasi-homogeneous domain to the dynamics of its automorphism group in Section 7.
6.1. A Kobayashi metric
In this section, the group $G$ is an $\operatorname{HTT}$ Lie group.
6.1.1. Chains.
Let $\Omega\subset\mathbf{Sb}(G)$ be a domain, not necessarily proper. We say that two points $x_{1},x_{2}\in\Omega$ are conjugate, denoted by $x_{1}\sim x_{2}$, if they belong to the same photon $\Lambda$ and are in the same connected component of $\Lambda\cap\Omega$.
Now let $x,y\in\Omega$ be any two points. An $N$-chain from $x$ to $y$ ($N\in\mathbb{N}$) is a sequence of $(N+1)$ elements $(x_{0}=x,\cdots,x_{N}=y)$ of $\Omega$ such that $x_{i}\sim x_{i+1}$ for all $10\leq i\leq N-$ 1. We denote by $\mathscr{C}_{x,y}(\Omega)$ (resp. $\mathscr{C}_{x,y}^{N}(\Omega)$) the set of all the chains (resp. $N$-chains) from $x$ to $y$ in $\Omega$.
In this section, we prove that there is a bound $\mathsf{n}(G)$ depending only on $G$ such that, locally, one can always join two points $x,y\in\mathbf{Sb}(G)$ by an $\mathsf{n}(G)$-chain.
Lemma 6.1.
There is an integer $n\geq 2$ only depending on $G$, such that for any two transverse points $p,q\in\mathbf{Sb}(B)$ and any diamond $D\subset\mathbf{Sb}(G)$ with endpoints $p$ and $q$, there is a sequence $(x_{0}=p,x_{1},\cdots,x_{n}=q)$ of $(n+1)$ elements of $\overline{D}$ such that $x_{i}$ and $x_{i+1}$ are on the same photon for all $10\leq i\leq n-$ 1.
  • Proof.
    Let $n$ be the smallest natural integer such that there exist $\ell_{1},...,\ell_{n}\in L^{0}$ such that $\sum_{k=1}^{n}\operatorname{Ad}(\ell_{k})\cdot v^{-}\in c^{0}$. The integer $n$ only depends on $G$. Define
    $x_{i}=\begin{cases}\ \ \ P&\text{ if }i=0;\\ \exp\big{(}\sum_{k=1}^{i}\operatorname{Ad}(\ell_{k})\cdot v^{-}\big{)}P&\text{if }1\leq i\leq n.\end{cases}$
    For $i\geq 1$, one has $x_{i}\in\mathbf{J}^{+}(x_{i-1})$, so by transitivity (Fact 3.6), one has $x_{i}\in\mathbf{J}^{+}(x_{0})$. In particular, one has $x_{n}\in\mathbf{I}^{+}(x_{0})$. Hence the set $D=\mathbf{I}^{+}(x_{0})\cap\mathbf{I}^{-}(x_{n})$ is a diamond. By construction, for any $10\leq i\leq n-$ 1, the points $x_{i}$ and $x_{i+1}$ are on the same photon, and all the $x_{i}$ are contained in $\overline{D}$.
[…]
Lemma 6.2.
There is an integer $\mathsf{n}(G)\geq 2$ depending only on $G$, such that for any diamond $D\subset\mathbf{Sb}(G)$, the set $\mathscr{C}^{\mathsf{n}(G)}_{x,y}(D)$ is nonempty for any $x,y\in D$.
  • Proof.
    We may assume that $D$ is of the form $D=\mathbf{D}(P,q)$, with $q\in\mathbf{I}^{+}(P)$; in particular, the domain $D$ is a convex subset of $\mathbb{A}$.
[…]
Remark 6.3.
The number $\mathsf{n}(G)$ found in the proof of Lemma 6.2 is not optimal. Indeed, for example when $G=\operatorname{SO}(n,2)$, the number $\mathsf{n}(G)=2$ works; see [CG24].
$q$$p$$z$$x$$y$
Figure 3. Example of path as constructed in the proof of Lemma 6.2, when $G=\operatorname{SO}(3,2)$. In this case, the number $\mathsf{n}(G)$ is not minimal; see Remark 6.3.
6.1.2. The pseudo-metric.
Let $x,y\in\Omega$ be two conjugate points. Let $\Lambda_{x,y}\in\mathscr{L}$ be the unique photon containing $x$ and $y$, and let $g\in G$ be such that $\Lambda_{x,y}=g\cdot\Lambda_{\mathsf{std}}$. We denote by $I_{x,y}$ the connected component of $\Omega\cap\Lambda_{x,y}$ containing $x$ and $y$. Recall the parametrization $\mathsf{i}_{g}:\mathbb{P}(\mathbb{R}^{2})\rightarrow\Lambda_{x,y}$ of (5.1). We define:
$\mathsf{k}_{\Omega}(x,y):=H_{\mathsf{i}_{g}^{-1}\left(I_{x,y}\right)}\left(\mathsf{i}_{g}^{-1}(x),\mathsf{i}_{g}^{-1}(y)\right).$
Recall that we denote by $H_{I}$ the Hilbert pseudo-metric of an interval $I$ of $\mathbb{P}(\mathbb{R}^{2})$ (see Section 2.1). Due to the $\operatorname{SL}_{2}(\mathbb{R})$-invariance of the cross ratio on $\mathbb{P}(\mathbb{R}^{2})$ and Lemma 5.4, the quantity $\mathsf{k}_{\Omega}(x,y)$ does not depend on the choice of $g\in G$ such that $\Lambda_{x,y}=g\cdot\Lambda_{\mathsf{std}}$.
Definition 6.4.
Given a domain $\Omega\subset\mathbf{Sb}(G)$, we define $K_{\Omega}:\Omega\times\Omega\rightarrow\mathbb{R}_{+}\cup\{+\infty\}$ by:
$\forall x,y\in\Omega,\quad K_{\Omega}(x,y)=\inf\left\{\sum_{i=0}^{N}\mathsf{k}_{\Omega}(x_{i},x_{i+1})\mid N\in\mathbb{N}^{*},\ \ (x_{0},\cdots x_{N})\in\mathscr{C}_{x,y}(\Omega)\right\}.$
For $x,y$ sufficiently close to each other, we can find a diamond included in $\Omega$ containing $x$ and $y$; see Remark 4.9.(1). In that case $\mathscr{C}_{x,y}(\Omega)$ (and even $\mathscr{C}^{\mathsf{n}(G)}_{x,y}(\Omega)$) is nonempty by Lemma 6.2. Since the relation “$x$ and $y$ can be joined by a chain” is an equivalence relation, the set $\mathscr{C}_{x,y}(\Omega)$ is never empty. Since $K_{\Omega}(a,b)$ is always finite whenever $a,b$ are conjugate, we know by Lemma 6.2 that the quantity $K_{\Omega}(x,y)$ is thus always finite as well. Thus $K_{\Omega}$ is actually a map $\Omega\times\Omega\rightarrow\mathbb{R}_{+}$.
We have the following standard naturality property:
Proposition 6.5.
Let $\Omega_{1}$ and $\Omega_{2}$ be two domains of $\mathbf{Sb}(G)$. Then:
  1. (1)
    If $\Omega_{1}\subset\Omega_{2}$, then for any $x,y\in\Omega_{1}$ one has $K_{\Omega_{2}}(x,y)\leq K_{\Omega_{1}}(x,y)$.
  2. (2)
    For any $g\in G$, for any $x,y\in\Omega_{1}$, one has $K_{g\cdot\Omega_{1}}(g\cdot x,g\cdot y)=K_{\Omega_{1}}(x,y)$. In particular, the metric $K_{\Omega_{1}}$ is $\mathsf{Aut}(\Omega_{1})$-invariant.
  • Proof.
[…]
Note that $K_{\Omega_{2}}$ and $K_{\Omega_{1}}$ do not need to be metrics in Proposition 6.5.
As one can concatenate and reverse the orientation of a chain, the map $K_{\Omega}$ is symmetric and satisfies the triangle inequality. It is thus a pseudo-metric, and we call it the Kobayashi pseudo-metric. In the next section, we investigate when $K_{\Omega}$ is a metric.
6.1.3. Kobayashi hyperbolicity
Let $\Omega\subset\mathbf{Sb}(G)$ be a domain. The domain $\Omega$ is said to be Kobayashi hyperbolic if $K_{\Omega}$ is a metric, that is, if $K_{\Omega}$ separates points.
Definition 6.6.
Let $\Omega\subset\mathbf{Sb}(G)$ be a Kobayashi-hyperbolic domain. Then the map $K_{\Omega}$ defined in Definition 6.4 is called the Kobayashi metric of $\Omega$.
Proposition 6.7.
Proper domains of $\mathbf{Sb}(G)$ are Kobayashi hyperbolic. Moreover, for such a domain $\Omega$, the metric $K_{\Omega}$ generates the standard topology.
  • Proof.
    Let us first show that a properly convex domain $C$ of $\mathbb{A}$ is Kobayashi hyperbolic, where $\mathbb{A}$ is endowed with its affine structure. Since $C$ is a properly convex domain of $\mathbb{A}$, it has a classical Hilbert metric $\mathsf{H}_{C}$ (see e.g. [Gol22]). By the definition of $\mathsf{H}_{C}$, if $a,b\in C$ are conjugate then one has
    (6.1) $\mathsf{H}_{C}(a,b)=\mathsf{k}_{C}(a,b).$
    Now let $x,y\in C$ distinct and $\gamma=(x_{0},\dots,x_{N})\in\mathscr{C}_{x,y}(\Omega)$ be a chain from $x$ to $y$. Since $\mathsf{H}_{C}$ satisfies the triangle inequality, one has:
    $\sum_{i=0}^{N-1}\mathsf{k}_{C}(x_{i},x_{i+1})=\sum_{i=0}^{N-1}\mathsf{H}_{C}(x_{i},x_{i+1})\geq\mathsf{H}_{C}(x,y).$
    This is true for all $\gamma\in\mathscr{C}_{x,y}(\Omega)$, so by taking the infinimum we get
    (6.2) $K_{C}(x,y)\geq\mathsf{H}_{C}(x,y)>0.$
    Therefore $K_{C}$ separates the points and $C$ is Kobayashi hyperbolic.
    Now let $\Omega$ be any domain. We may assume that $\Omega$ is proper in $\mathbb{A}$ (Remark 4.2). Let $C$ be any properly convex domain of $\mathbb{A}$ containing $\Omega$ (for instance the convex hull of $\Omega$ in $\mathbb{A}$). Then, by Proposition 6.5:
    (6.3) $K_{C}(x,y)\leq K_{\Omega}(x,y)\quad\forall x,y\in\Omega.$
    Since $K_{C}$ separates the points, so does $K_{\Omega}$.
    Let us show that $K_{\Omega}$ generates the standard topology. Let $\mathscr{T}_{\mathsf{std}}$ be the standard topology on $\Omega$ and $\mathscr{T}$ the topology induced by $K_{\Omega}$. We will prove that $\mathscr{T}_{\mathsf{std}}=\mathscr{T}$. By Equations (6.2) and (6.3) and the fact that the Hilbert metric generates the standard topology on properly convex domains, one has $\mathscr{T}_{\mathsf{std}}\subset\mathscr{T}$. To prove the reverse inclusion, one need to show that $K_{\Omega}$ is continuous with respect to the standard topology. By the inequality
    $|K_{\Omega}(x_{0},y_{0})-K_{\Omega}(x,y)|\leq K_{\Omega}(x_{0},x)+K_{\Omega}(y_{0},y)\quad\forall x_{0},y_{0},x,y\in\Omega,$
    one only needs to show that for any $x_{0}\in\Omega$ the map $x\mapsto K_{\Omega}(x_{0},x)$ is continuous at $x_{0}$. For this we see $\mathbb{A}$ as an Euclidean space and we denote by ||.||||.||| | . | | the associated Euclidean norm. Up to dilating at $x_{0}$, we may assume that the Euclidean ball $B$ of center $x_{0}$ and of radius $1$ is contained in $\Omega$. Let $D$ be a diamond containing $x_{0}$ and contained in $B$, with diameter $\delta\in]0,1[$ for the Euclidean norm (Remark 4.9.(1)). Let $N:=\mathsf{n}(G)$ given by Lemma 6.2. There exists an $N$-chain $(x_{0},x_{1},\cdots,\ x_{N}=x)$ contained in $D\subset B$. Then, by Proposition 6.5 and Equation (6.1), one has
    $K_{\Omega}(x,x_{0})$ $\leq K_{B}(x_{0},x)\leq\sum_{k=0}^{N-1}K_{B}(x_{i},x_{i+1})=\sum_{k=0}^{N-1}\mathsf{H}_{B}(x_{i},x_{i+1})$
    $\leq\sum_{k=0}^{N-1}\mathsf{H}_{B}(x_{0},x_{i})+\mathsf{H}_{B}(x_{0},x_{i+1})$
    $=\sum_{k=1}^{N-1}\log\frac{1+||x_{1}-x_{0}||}{1-||x_{i}-x_{0}||}+\log\frac{1+||x_{i+1}-x_{0}||}{1-||x-x_{0}||}\leq\sum_{k=1}^{N-1}2\log\frac{1+\delta}{1-\delta}\ \underset{\delta\rightarrow 0}{\longrightarrow}0$
[…]
6.1.4. Kobayashi length.
In this section, we recall some definitions and fix some notation. Let $\Omega$ be a Kobayashi hyperbolic domain, contained in $\mathbb{A}$. For a continuous path $\gamma:[0,1]\rightarrow\Omega$, we define the Kobayashi length or $K_{\Omega}$-length of $\gamma$ in the the usual way, as
$\mathsf{len}_{\Omega}(\gamma)=\sup\sum_{i=0}^{N}K_{\Omega}(\gamma(t_{i}),\gamma(t_{i+1})),$
where the supremum is taken over all finite subdivisions of $\gamma$.
Let $x,y\in\Omega$ be two conjugate points, and let $\Lambda_{x,y}$ be the unique photon containing $x$ and $y$. We denote by $[x,y]$ the closure of the only connected component of $\Lambda_{x,y}\smallsetminus\{x,y\}$ that is contained in $\Omega$. By Lemma 5.5, it is an affine segment in $\mathbb{A}$. This segment can be parametrized by
$[t_{1},t_{2}]\longrightarrow[x,y];\ t\longmapsto\mathsf{i}_{g}([1:t]),$
where $g\in G$ is such that $\Lambda_{x,y}=g\cdot\Lambda_{\mathsf{std}}$ and $x=\mathsf{i}_{g}([1:t_{1}])$ and $y=\mathsf{i}_{g}([1:t_{2}])$. This parametrization depends on the choice of $g\in G$.
Now, let $x,y\in\Omega$ be any two points. Any element of $u=(x_{0},\cdots,x_{N})\in\mathscr{C}_{x,y}(\Omega)$ gives rise to a continuous path $\gamma$ from $x$ to $y$, defined as the concatenation of all the segments $[x_{0},x_{1}],\cdots,[x_{N-1},x_{N}]$ in this order, endowed with a parametrization as described above. This path is uniquely defined by $u$ up to reparametrization. The $K_{\Omega}$-length $\mathsf{len}_{\Omega}(\gamma)$ of $\gamma$ does not depend on the choice of parametrization of the $[x_{i},x_{i+1}]$ for $0\leq i\leq N$. This defines a unique $K_{\Omega}$-length for the chain $(x_{0},\cdots,x_{N})$.
In the rest of the paper, we will identify a chain with the unique (up to parametrization) path it defines by the process described above. In particular, this will allow us to consider the $K_{\Omega}$-length of a chain.
6.2. Comparison with the Caratheodory metrics
The aim of this section is to prove Proposition 1.9. The results we prove are actually stronger; see Proposition 6.8 and Corollary 6.13.
6.2.1. The Caratheodory metric
Let $G$ be a real semisimple Lie group and $\Theta\subset\operatorname{\Delta}$ be a subset of the simple roots. Let $(V,\rho)$ be a $\Theta$-proximal finite-dimensional real irreducible linear representation of $G$ (see Section 2.4). Let $\iota:\mathscr{F}_{\Theta}\hookrightarrow\mathbb{P}(V)$ and $\iota^{*}:\mathscr{F}_{\Theta}^{-}\hookrightarrow\mathbb{P}(V^{*})$ be the two embeddings induced by $(V,\rho)$, see Proposition 2.1. Given $x,y\in\mathscr{F}_{\Theta}$ and $\xi,\eta\in\mathscr{F}_{\Theta}^{-}$, let
$\nu_{x}\in\iota(x)\smallsetminus\{0\};\quad\nu_{y}\in\iota(y)\smallsetminus\{0\};\quad f_{\xi}\in\iota^{*}(\xi)\smallsetminus\{0\};\quad f_{\eta}\in\iota^{*}(\eta)\smallsetminus\{0\}.$
We define the cross ratio of $\xi,x,y,\eta$ relative to $(V,\rho)$ as follows:
(6.4) $\left[\xi:x:y:\eta\right]_{\rho}:=\frac{f_{\xi}(\nu_{x})f_{\eta}(\nu_{y})}{f_{\xi}(\nu_{y})f_{\eta}(\nu_{x})}.$
This quantity does not depend on the choice of representatives $\nu_{x},\nu_{y},f_{\xi},f_{\eta}$. In [Zim18a], A. Zimmer introduces the following map $C_{\Omega}$ associated with a domain $\Omega\subset\mathscr{F}_{\Theta}$:
$C_{\Omega}:\begin{cases}\Omega\times\Omega&\longrightarrow\quad\quad\mathbb{R}_{+}\\ (x,y)&\longmapsto\sup_{\xi,\eta\in\Omega^{*}}\log\big{|}\left[\xi:x:y:\eta\right]_{\rho}\big{|}.\end{cases}$
By [Zim18a, Thm 5.2 and 9.1], the map $C_{\Omega}$ is an $\mathsf{Aut}(\Omega)$-invariant metric generating the standard topology, as soon as $\Omega$ is a proper domain of $\mathscr{F}_{\Theta}$. Whenever this is the case, we will say that $C_{\Omega}$ is the Caratheodory metric on $\Omega$ induced by $(V,\rho)$.
6.2.2. Comparison for Shilov boundaries.
The goal of this section is to prove Proposition 6.8 below. Together with Corollary 6.13 in the next section, it will imply Proposition 1.9.
Proposition 6.8.
Let $G$ be a simple Lie group of Hermitian type of tube type and let $(V,\rho)$ be a finite-dimensional real irreducible linear representation of $G$ with highest weight $\chi=N\omega_{\alpha_{r}}$ ($N\in\mathbb{N}_{>0}$). Let $\Omega\subset\mathbf{Sb}(G)$ be a proper dually convex domain, and let $C_{\Omega}$ be the Caratheodory metric on $\Omega$ induced by $(V,\rho)$. Then for any two conjugate points $x,y\in\Omega$, one has
$\mathsf{k}_{\Omega}(x,y)=\frac{1}{N}C_{\Omega}(x,y).$
In particular,
  1. (1)
    one has $K_{\Omega}\geq N^{-1}C_{\Omega}$;
  2. (2)
    the metric $K_{\Omega}$ is a length metric, and given two conjugate points $x,y\in\Omega$, the $K_{\Omega}$-length of the $1$-chain $(x,y)$ is equal to $\mathsf{k}_{\Omega}(x,y)=K_{\Omega}(x,y)$.
Let us fix once and for all some notation for the rest of this section. Let $G$ be an $\operatorname{HTT}$ Lie group. Let $(V,\rho)$ be a finite-dimensional real irreducible linear representation of $G$ with highest weight $\chi=N\omega_{r}$ for some $N\in\mathbb{N}_{>0}$. We let $\iota:\mathbf{Sb}(G)\hookrightarrow\mathbb{P}(V)$ and $\iota^{*}:\mathbf{Sb}(G)\hookrightarrow\mathbb{P}(V^{*})$ be the two embeddings induced by $\rho$ from Proposition 2.1, and we fix a vector $e_{1}\in V^{\chi}\smallsetminus\{0\}$.
The key point of the proof of Proposition 6.8 is the comparison of two cross ratios: the one defined in Equation (6.4) and the one given in the definition of $\mathsf{k}_{\Omega}$. This is done in Lemma 6.11 below. The argument uses the intersection polynomials introduced in Section 5.2. We have seen that the real roots of an intersection polynomial $\mathsf{Q}_{\xi}^{\rho}$ for $\xi\in\mathscr{X}$ are geometrically described by the intersection points of $\iota(\Lambda_{\mathsf{std}})$ with $\iota^{*}(\xi)$, so that they have at most one real root by Lemma 5.5. But our argument will rely on the fact that they have only one complex root, so that they are split (see Corollary 6.10). The complex roots of $\mathsf{Q}_{\xi}^{\rho}$ will be geometrically described as the intersection points of the complexification of $\iota^{*}(\xi)$ and a set, denoted $\mathscr{P}_{N}$, that plays the role of the complexification of $\iota(\Lambda_{\mathsf{std}})$. We describe this intersection in Lemma 6.9 below and actually prove that it is still a singleton. To this end, we work in a complexification of the representation $(V,\rho)$.
Let $V^{\mathbb{C}}:=V\otimes\mathbb{C}$ be the complexification of $V$. Recall the notation of Section 2.1.
For any $\xi\in\mathbf{Sb}(G)$ and any lift $f\in V^{*}\smallsetminus\{0\}$ of $\iota^{*}(\xi)$, the map $f$ extends uniquely to a linear form $f^{\mathbb{C}}:V^{\mathbb{C}}\rightarrow\mathbb{C}$. We denote by $\iota^{*}(\xi)^{\mathbb{C}}$ the class $[f^{\mathbb{C}}]_{c}$ of $f^{\mathbb{C}}$ in $\mathbb{P}_{c}\left((V^{\mathbb{C}})^{*}\right)$. This definition does not depend on the choice of the lift $f$ of $\iota^{*}(\xi)$ in $V^{*}\smallsetminus\{0\}$. As for the real case, we identify $\mathbb{P}_{c}\left((V^{\mathbb{C}})^{*}\right)$ with the set of projective hyperplanes of $\mathbb{P}_{c}(V^{\mathbb{C}})$.
Similarly, for any $g\in G$, the operator $\rho(g)$ uniquely extends to an automorphism of $V^{\mathbb{C}}$. We will still denote by $\rho(g)$ this extension.
The map $\tau:\operatorname{SL}_{2}(\mathbb{R})\rightarrow G$ induces a group homomorphism $\operatorname{SL}_{2}(\mathbb{R})\rightarrow G\rightarrow\operatorname{GL}(V)$ with kernel included in {±$\{\pm\operatorname{id}\}$. It extends to a homomorphism $\tau_{\mathbb{C}}:\operatorname{SL}_{2}(\mathbb{C})\rightarrow\operatorname{GL}(V^{\mathbb{C}})$ (with kernel {±$\{\pm\operatorname{id}\}$), such that
$\tau_{\mathbb{C}}\left(e^{\operatorname{E}}\right)=e^{\rho_{*}(v^{+})},\quad\tau_{\mathbb{C}}\left(e^{\operatorname{F}}\right)=e^{\rho_{*}(v^{-})},\quad\tau_{\mathbb{C}}\left(e^{\operatorname{H}}\right)=e^{\rho_{*}(h_{r})}.$
The stabilizer of $e_{1}$ in $\operatorname{SL}_{2}(\mathbb{C})$ is the standard Borel subgroup $P_{1}^{\mathbb{C}}$ of $\operatorname{SL}_{2}(\mathbb{C})$. Hence the map $\tau_{\mathbb{C}}$ induces a $\tau_{\mathbb{C}}$-equivariant embedding $\mathsf{F}:\mathbb{P}_{c}(\mathbb{C}^{2})\hookrightarrow\mathbb{P}_{c}(V^{\mathbb{C}})$. The image of $\mathsf{F}$ is denoted by $\mathscr{P}_{N}$. Explicitly, the set $\mathscr{P}_{N}$ is the closure in $\mathbb{P}_{c}(V^{\mathbb{C}})$ of
$\big{\{}\mathsf{F}\left([1:z]_{c}\right)=e^{z\rho_{*}(v^{-})}\cdot[e_{1}]_{c}\ \ \big{|}\ z\in\mathbb{C}\big{\}}.$
Lemma 6.9.
  1. (1)
    One has $\rho(U^{+})\cdot\mathscr{P}_{N}=\mathscr{P}_{N}$.
  2. (2)
    Let $\xi\in\mathscr{X}$. Then the set $\mathscr{P}_{N}\cap\iota^{*}(\xi)^{\mathbb{C}}$ is a singleton.
  • Proof.
    Let us first prove (1). Note that this is not just a consequence of Lemma 5.2, although it has a similar proof. As in the proof of Lemma 3.8, for all $Y\in\mathfrak{u}^{+}$, using Lemma 3.2, we can find $\lambda\in\mathbb{R}$ such that for all $z\in\mathbb{C}\smallsetminus\{-\lambda^{-1}\}$ (with $-\lambda^{-1}=\infty$ if $\lambda=0$), one has
    $\rho\left(\exp(Y)\right)\cdot\mathsf{F}\left([1:z]_{c}\right)=\mathsf{F}\left(\left[1:\frac{z}{1+z\lambda}\right]_{c}\right)\in\mathscr{P}_{N}.$
    Hence $\rho\left(\exp(Y)\right)\cdot\mathsf{F}\left([1:z]_{c}\right)\in\mathscr{P}_{N}$ for all $z\in\mathbb{C}\smallsetminus\{-\lambda^{-1}\}$. Taking the closure, we get $\rho\left(\exp(Y)\right)\cdot\mathscr{P}_{N}\subset\mathscr{P}_{N}$. The converse inclusion also holds by the same argument applied to $-Y$ instead of $Y$. Therefore $\rho\left(\exp(Y)\right)\cdot\mathscr{P}_{N}=\mathscr{P}_{N}$.
Now let us prove (2).
Step 1. Let us first prove (2) for $\xi=P^{-}\in\mathscr{X}$. Let $f_{0}\in V^{*}\smallsetminus\{0\}$ be any lift of $\iota^{*}(P^{-})$ and let $x\in\iota^{*}(P^{-})^{\mathbb{C}}\cap\mathscr{P}_{N}$.
There exists $(z_{n})\in\mathbb{C}^{\mathbb{N}}$ be such that $\mathsf{F}\left(e^{z_{n}\operatorname{F}}\cdot[1:0]_{c}\right)\rightarrow x$ as $n\rightarrow+\infty$. Since
$f_{0}^{\mathbb{C}}\big{(}e^{z\rho_{*}(v^{-})}e_{1}\big{)}=f_{0}(e_{1})+zf_{0}\left(\rho_{*}(v^{-})e_{1}\right)+\cdots+\frac{z^{N}}{N!}f_{0}\left(\rho_{*}(v^{-})^{N}e_{1}\right)=f_{0}(e_{1})\neq 0$
for all $z\in\mathbb{C}$, we must have $|z_{n}|\rightarrow+\infty$. Then $e^{z_{n}\operatorname{F}}\cdot[1:0]_{c}\rightarrow[0:1]_{c}$. Hence $x$ has to be equal to $\mathsf{F}\left([0:1]_{c}\right)$.
Step 2. Now let $\xi\in\mathscr{X}$ be any point, and let $g\in G$ be such that $\xi=g^{-1}P^{-}$. Since $\Lambda_{\mathsf{std}}\not\subset\operatorname{Z}_{\xi}$, there exists $t\in\mathbb{R}$ such that $g^{-1}\exp(tv^{-})P\in\mathbb{A}$. Since $\exp(tv^{-})$ preserves $P^{-}$, one has $g^{-1}\exp(tv^{-})P^{-}=\xi$. Hence, up to replacing $g$ with $g^{-1}\exp(tv^{-})$, we may assume that $gP\in\mathbb{A}$. Then, by (2.4), we can write $g=h\exp(Y)$ with $Y\in\mathfrak{u}^{+}$ and $h\in P^{-}$. By point (1) of the lemma, one has $\rho\left(\exp(Y)\right)\cdot\mathscr{P}_{N}=\mathscr{P}_{N}$. On the other hand, since $\rho(h^{-1})$ preserves $\iota^{*}(P^{-})$, its $\mathbb{C}$-extension preserves $\iota^{*}(P^{-})^{\mathbb{C}}$. Thus $\rho(h^{-1})\cdot\iota^{*}(P^{-})^{\mathbb{C}}=\iota^{*}(P^{-})^{\mathbb{C}}$. This gives
$\mathscr{P}_{N}\cap\iota^{*}(\xi)^{\mathbb{C}}$ $=\mathscr{P}_{N}\cap\left(\rho\left(\exp(-Y)\right)\rho\left(h^{-1}\right)\cdot\iota^{*}(P^{-})^{\mathbb{C}}\right)$
$=\rho\left(\exp(-Y)\right)\left(\rho(\exp(Y))\mathscr{P}_{N}\cap\iota^{*}(P^{-})^{\mathbb{C}}\right)$
$=\rho\left(\exp(-Y)\right)\left(\mathscr{P}_{N}\cap\iota^{*}(P^{-})^{\mathbb{C}}\right),$
[…]
Lemma 6.9.(2) above admits the following corollary:
Corollary 6.10.
Let $\xi\in\mathscr{X}$. Then the intersection polynomial $\mathsf{Q}_{\xi}^{\rho}$ of $\xi$ is split. If moreover $\Lambda_{\mathsf{std}}\cap\operatorname{Z}_{\xi}\subset\mathbb{A}$, then the unique complex root of $\mathsf{Q}_{\xi}^{\rho}$ is equal to the unique $t\in\mathbb{R}$ satisfying $\mathsf{i}([1:t])\in\operatorname{Z}_{\xi}$.
  • Proof.
    With the notation of Section 5.2, let $f\in V^{*}$ be the unique lift of $\xi$ such that $f(\rho_{*}(v^{-})^{n(\xi)}e_{1})=n(\xi)!$. For all $z\in\mathbb{C}$, one has:
    $\mathsf{Q}_{\xi}^{\rho}(z)=f(e_{1})+zf(\rho_{*}(v^{-})e_{1})+\cdots+z^{n(\xi)}=f^{\mathbb{C}}\big{(}e^{z\rho_{*}(v^{-})}\cdot e_{1}\big{)}$
    Hence one has:
    (6.5) $\mathsf{Q}_{\xi}^{\rho}(z)=0\ \Longleftrightarrow\ f^{\mathbb{C}}\big{(}e^{z\rho_{*}(v^{-})}\cdot e_{1}\big{)}=0\ \Longleftrightarrow\ \mathsf{F}\left([1:z]_{c}\right)\in\mathscr{P}_{N}\cap\iota^{*}(\xi)^{\mathbb{C}}.$
    By Lemma 6.9.(2), the intersection $\mathscr{P}_{N}\cap\iota^{*}(\xi)^{\mathbb{C}}$ is a singleton, so the injectivity of $\mathsf{F}$ and the equivalence of (6.5) above give that $\mathsf{Q}_{\xi}^{\rho}$ has only one complex root. Therefore, the polynomial $\mathsf{Q}_{\xi}^{\rho}$ is split.
[…]
Recall that, by Lemma 5.5, given a point $\xi\in\mathbf{Sb}(G)\smallsetminus\Lambda_{\mathsf{std}}=\mathscr{X}$, the set $\operatorname{Z}_{\xi}\cap\Lambda_{\mathsf{std}}$ is a singleton. We can then define a $G$-equivariant projection $\operatorname{pr}_{\mathsf{std}}:\mathbf{Sb}(G)\rightarrow\Lambda_{\mathsf{std}}$ by setting $\operatorname{pr}_{\mathsf{std}}(\xi):=\xi$ if $\xi\in\Lambda_{\mathsf{std}}$, and $\operatorname{pr}_{\mathsf{std}}(\xi):=p$ where $\{p\}=\operatorname{Z}_{\xi}\cap\Lambda_{\mathsf{std}}$ if $\xi\in\mathscr{X}$. Using Lemma 6.9 and Corollary 6.10, we now establish a comparison between two cross ratios, involving the projection $\operatorname{pr}_{\mathsf{std}}$:
Lemma 6.11.
Let $\xi_{1},\xi_{2}\in\mathbf{Sb}(G)$, and for $i\in\{1,2\}$, let $b_{i}\in\mathbb{P}(\mathbb{R}^{2})$ be such that $\operatorname{pr}_{\mathsf{std}}(\xi_{i})=\mathsf{i}(b_{i})$. Let $a_{1},a_{2}\in\mathbb{P}(\mathbb{R}^{2})$ be such that $b_{1},a_{1},a_{2},b_{2}\in\mathbb{P}(\mathbb{R}^{2})$ are aligned in this order. Then one has
$\log\left|[\xi_{1}:\mathsf{i}(a_{1}):\mathsf{i}(a_{2}):\xi_{2}]_{\rho}\right|=N\log(b_{1}:a_{1}:a_{2}:b_{2}).$
Given two points $x,y\in\Lambda_{\mathsf{std}}$, Lemma 6.11 expresses that the cross ratio $[\xi_{1},x,y,\xi_{2}]_{\rho}$, where $\xi_{1},\xi_{2}\in\mathbf{Sb}(G)$, depends only on the projections of $\xi_{1}$ and $\xi_{2}$ to $\Lambda_{\mathsf{std}}$.
Remark 6.12.
After this paper was completed, we learnt that Beyrer–Guichard–Labourie–Pozzetti–Wienhard [BGL+] prove a similar property for flag manifolds with a $\Theta$-positive structure. Other versions are established in [Zim15, Lem. 10.4] for Grassmannians $\operatorname{Gr}_{p}(\mathbb{R}^{n})$ and in [LZ17, Lem. 2.9] for the full flag manifold of $\operatorname{SL}(n,\mathbb{R})$.
  • Proof of Lemma 6.11.
    Let us set $x:=P=\mathsf{i}(a_{1})$, $y:=\mathsf{i}(a_{2})$, $p_{1}=\mathsf{i}(b_{1})$ and $p_{2}=\mathsf{i}(b_{2})$. We may assume that $a_{1}=[1:0]$ and $a_{2}=[1:1]$, and that there exist $t_{1},t_{2}\in\mathbb{R}$ such that $b_{i}=[1:t_{i}]$ for $i\in\{1,2\}$, and that $t_{1}<0<1<t_{2}$, the lemma then following by a continuity argument. Then the four distinct points $p_{1},x,y,p_{2}$ are aligned on $\Lambda_{\mathsf{std}}\cap\mathbb{A}$ in this order. Note that $\iota(x)=[e_{1}]$. Let $e_{2}:=e_{1}+\rho_{*}(v^{-})e_{1}+\cdots+(1/N!)\rho_{*}(v^{-})^{N}e_{1}$. Then, by (5.3), one has $\iota(y)=[e_{2}]$.
    Recall the open set $\mathscr{A}_{\rho}$ of (5.5). By Lemma 5.8, the set $\mathscr{A}_{\rho}$ is a dense open subset of $\mathbf{Sb}(G)$. Hence for $i\in\{1,2\}$, we can find a sequence $(\xi_{i,n})\in\mathscr{A}_{\rho}^{\mathbb{N}}$ such that $\xi_{i,n}\rightarrow\xi_{i}$. For all $n\in\mathbb{N}$, let $p_{i,n}:=\operatorname{pr}_{\mathsf{std}}(\xi_{i,n})$. Then, by continuity of $\operatorname{pr}_{\mathsf{std}}$, one has $p_{i,n}\rightarrow p_{i}\in\mathbb{A}$, so up to extracting we may assume that $p_{i,n}\in\mathbb{A}$ for all $n\in\mathbb{N}$.
    Let $f_{i}\in V^{*}\smallsetminus\{0\}$ (resp. $f_{i,n}\in V^{*}\smallsetminus\{0\}$) be a lift of $\iota^{*}(\xi_{i})$ (resp.  $\iota^{*}(\xi_{i,n})$). For every $n\in\mathbb{N}$ we choose $f_{i,n}$ such that $f_{i,n}(\rho_{*}(v^{-})^{N}e_{1})=N!$. For any $n\in\mathbb{N}$ and $i\in\{1,2\}$, the intersection polynomial
    (6.6) $\mathsf{Q}_{i,n}(z):=\mathsf{Q}_{\xi_{i,n}}^{\rho}(z)=(f_{i,n})^{\mathbb{C}}(e^{z\rho_{*}(v^{-})}e_{1})=f_{i,n}(e_{1})+zf_{i,n}(\rho_{*}(v^{-})e_{1})+\cdots+z^{N}$
    is nonzero, so there exists $t\in\mathbb{R}$ such that $\mathsf{Q}_{i,n}(t)=f_{i,n}(e^{t\rho_{*}v^{-}}e_{1})\neq 0$. This implies in particular that $\Lambda_{\mathsf{std}}\not\subset\operatorname{Z}_{\xi_{i,n}}$, i.e. $\xi\in\mathscr{X}$. By Corollary 6.10, the polynomial $\mathsf{Q}_{i,n}$ is thus split. But we also know that $\Lambda_{\mathsf{std}}\cap\operatorname{Z}_{\xi_{i,n}}=\{p_{i,n}\}$ is contained in $\mathbb{A}$. Hence by the “moreover” part of Corollary 6.10, and since $n(\xi_{i,n})=N$, the polynomial $\mathsf{Q}_{i,n}$ can be written $\mathsf{Q}_{i,n}(z)=(z-t_{i,n})^{N}$, with $t_{i,n}\in\mathbb{R}$ satisfying $\mathsf{i}([1:t_{i,n}])=p_{i,n}$. Since $(p_{i,n})$ converges to $p_{i}$, the sequence $(t_{i,n})$ converges to $t_{i}$. One then has:
    $\log(s_{1}:[1:0]:[1:1]:s_{2})$ $=\log\left|\frac{t_{1}\cdot(t_{2}-1)}{t_{2}\cdot(t_{1}-1)}\right|$
    $=\lim_{n\rightarrow+\infty}\log\left|\frac{t_{1,n}\cdot(t_{2,n}-1)}{t_{2,n}\cdot(t_{1,n}-1)}\right|$
    $=\frac{1}{N}\lim_{n\rightarrow+\infty}\log\left|\frac{\mathsf{Q}_{1,n}(0)\mathsf{Q}_{2,n}(1)}{\mathsf{Q}_{2,n}(0)\mathsf{Q}_{1,n}(1)}\right|$
    $=\frac{1}{N}\lim_{n\rightarrow+\infty}\log\left|\frac{f_{1,n}(e_{1})f_{2,n}(e_{2})}{f_{2,n}(e_{1})f_{2,n}(e_{2})}\right|$
    $=\frac{1}{N}\log\left|\frac{f_{1}(e_{1})f_{2}(e_{2})}{f_{2}(e_{1})f_{2}(e_{2})}\right|=\frac{1}{N}\log\big{|}\left[\xi_{1}:x:y:\xi_{2}\right]_{\rho}\big{|}.\quad\hfill\square$
  • Proof of Proposition 6.8.
    Up to translating $\Omega$ by an element of $G$, one may assume that $x,y\in\Lambda_{\mathsf{std}}$. Then there exist $a_{1},a_{2}\in\mathbb{P}(\mathbb{R}^{2})$ such that $x=\mathsf{i}(a_{1})$ and $y=\mathsf{i}(a_{2})$.
    Recall that we denote by $I_{x,y}$ the connected component of $\Lambda_{\mathsf{std}}\cap\Omega$ containing $x$ and $y$. Let $p_{1},p_{2}\in\partial\Omega$ be the endpoints of $I_{x,y}$, such that $p_{1},x,y,p_{2}$ are aligned on $\Lambda_{\mathsf{std}}$ in this order. Then there exist $b_{1},b_{2}\in\mathbb{P}(\mathbb{R}^{2})$ such that $b_{1},a_{1},a_{2},b_{2}$ are aligned in this order and $p_{1}=\mathsf{i}(b_{1})$, and $p_{2}=\mathsf{i}(b_{2})$.
    By dual convexity, for $i\in\{1,2\}$ there exists $\xi_{i}\in\Omega^{*}$ such that $p_{i}\in\operatorname{Z}_{\xi_{i}}$. Then, by Lemma 6.11, one has
    $\mathsf{k}_{\Omega}(x,y)=\log\left|(b_{1}:a_{1}:a_{2}:b_{2})\right|=\frac{1}{N}\log\big{|}\left[\xi_{1}:x:y:\xi_{2}\right]_{\rho}\big{|}.$
    By the definition of $C_{\Omega}$, this implies that $\mathsf{k}_{\Omega}(x,y)\leq N^{-1}C_{\Omega}(x,y)$.
    For the converse inequality, let $\eta_{1},\eta_{2}\in\Omega^{*}$ be such that $C_{\Omega}(x,y)=\log\left|[\eta_{1}:x:y:\eta_{2}]_{\rho}\right|$. For $i\in\{1,2\}$, let $b_{i}^{\prime}\in\mathbb{P}(\mathbb{R}^{2})$ be such that $\mathsf{i}(b_{i}^{\prime})=\operatorname{pr}_{\mathsf{std}}(\eta_{i})$. Then, again by Lemma 6.11:
    $\big{|}\log\left|(b_{1}^{\prime}:a_{1}:a_{2}:b_{2}^{\prime})\right|\big{|}=\frac{1}{N}\log\left|[\eta_{1}:x:y:\eta_{2}]_{\rho}\right|.$
    Since $\eta_{1},\eta_{2}\in\Omega^{*}$, the two points $\mathsf{i}(b_{1}^{\prime}),\mathsf{i}(b_{2}^{\prime})$ are not contained in $I_{x,y}$. Thus one has
    $\big{|}\log\left|(b_{1}^{\prime}:a_{1}:a_{2}:b_{2}^{\prime})\right|\big{|}\leq\log\left|\left(b_{1}:a_{1}:a_{2}:b_{2}\right)\right|=\mathsf{k}_{\Omega}(x,y).$
    Hence one has $N^{-1}C_{\Omega}(x,y)\leq\mathsf{k}_{\Omega}(x,y)$. We have proved that $N^{-1}C_{\Omega}(x,y)=\mathsf{k}_{\Omega}(x,y)$.
    Now let us prove that $K_{\Omega}\geq N^{-1}C_{\Omega}$. Let $x,y\in\Omega$ be any two points, and let $(x_{0},\cdots,x_{M})\in\mathscr{C}_{x,y}(\Omega)$. Then one has
    (6.7) $\begin{split}\sum_{i}\mathsf{k}_{\Omega}(x_{i},x_{i+1})=\frac{1}{N}\sum_{i}C_{\Omega}(x_{i},x_{i+1})&\geq\sum_{i}\frac{1}{N}\sup_{\xi_{1}^{i},\xi_{2}^{i}\in\Omega^{*}}\log[\xi_{1}^{i}:x_{i}:x_{i+1}:\xi_{2}^{i}]_{\rho}\\ &\geq\frac{1}{N}\sup_{\xi_{1},\xi_{2}\in\Omega^{*}}\sum_{i}\log[\xi_{1}:x_{i}:x_{i+1}:\xi_{2}]_{\rho}\\ &\geq\frac{1}{N}\sup_{\xi_{1},\xi_{2}\in\Omega^{*}}\log[\xi_{1}:x:y:\xi_{2}]_{\rho}\\ &=\frac{1}{N}C_{\Omega}(x,y).\end{split}$
    Since this is true for all $(x_{0},\cdots,x_{M})\in\mathscr{C}_{x,y}(\Omega)$, by taking the infimum we get the inequality $K_{\Omega}(x,y)\geq N^{-1}C_{\Omega}(x,y)$.
    Now let us show that $K_{\Omega}$ is a length metric. In (6.7), take $x$ and $y$ to be two conjugate points. The fact that $C_{\Omega}(x,y)=N\mathsf{k}_{\Omega}(x,y)$ and Equation (6.7) imply that the segment $[x,y]$ has $K_{\Omega}$-length $\mathsf{k}_{\Omega}(x,y)$. Hence the $K_{\Omega}$-length of $\gamma=(x_{0},\cdots,x_{M})\in\mathscr{C}_{x,y}(\Omega)$ is
    (6.8) $\mathsf{len}_{\Omega}(\gamma)=\sum_{i}K_{\Omega}(x_{i},x_{i+1}).$
    Then one has $K_{\Omega}(x,y)=\inf\left\{\mathsf{len}_{\Omega}(\gamma)\mid\gamma\in\mathscr{C}_{x,y}(\Omega)\right\}$.
[…]
6.2.3. Properness of $K_{\Omega}$.
Let us fix a real finite-dimensional real irreducible linear representation $(V,\rho)$ of $G$ with highest weight $N\omega_{r}$ for some $N\in\mathbb{N}_{>0}$. Let $\Omega$ be a proper domain of $\mathbf{Sb}(G)$, and let $C_{\Omega}$ be the Caratheodory metric on $\Omega$ induced by $(V,\rho)$. In [Zim18a, Thm 9.1], Zimmer proves that the following three assertions are equivalent:
  1. (1)
    $\Omega$ is dually convex;
  2. (2)
    $C_{\Omega}$ is a proper metric;
  3. (3)
    $C_{\Omega}$ is a complete metric.
The equivalence $(1)\Leftrightarrow(3)$ is stated in [Zim18a, Thm 9.11], and the equivalence $(1)\Leftrightarrow(2)$ is a consequence of its proof.
By Proposition 6.8, one has $K_{\Omega}\geq N^{-1}C_{\Omega}$. Thus the Kobayashi metric $K_{\Omega}$ is also proper:
Corollary 6.13.
If $\Omega\subset\mathbf{Sb}(G)$ is a proper dually convex domain, then $K_{\Omega}$ is a proper metric. In particular, if $\Omega\subset\mathbf{Sb}(G)$ is a proper quasi-homogeneous domain, then $K_{\Omega}$ is a proper metric.
Proposition 6.8 and Corollary 6.13 imply in particular that $K_{\Omega}$ is geodesic as soon as $\Omega$ is a proper quasi-homogeneous (and even just dually convex) domain, although we will not use this fact.
Remark 6.14.
The only consequence of Proposition 6.8 and Corollary 6.13 we will use in the proof of Theorem 1.4 is the fact that $K_{\Omega}$ is proper as soon as $\Omega$ is quasi-homogeneous, and that the $K_{\Omega}$-length of a $1$-chain $(x,y)$ is in this case equal to $\mathsf{k}_{\Omega}(x,y)$. Other approaches would have led to this conclusion: for instance, the theory of Euclidean Jordan algebras, or looking at explicit proximal representations of $G$ for each value of $\mathfrak{g}$ in Table 1. However, the approach we choose makes it possible to compare the Kobayashi metric with the already existing Caratheodory metrics, and is generalizable to other flag manifolds in which Kobayashi constructions have already been made (see Remark 1.10).
7. The dynamics at the boundary
This section is divided into two parts. Section 7.1 is devoted to the definition and the analysis of two families of points of the boundary of a proper domain in $\mathbf{Sb}(G)$, those of visual and $\mathscr{R}$-extremal points of the domain. In Section 7.2, we prove the existence of $\mathscr{R}$-extremal points satisfying a particular geometric property, namely strongly $\mathscr{R}$-extremal points.
7.1. Visual points and $\mathscr{R}$-extremal points
In this section, following [LZ19], we investigate the relation between the structure of the boundary of a proper quasi-homogeneous domain and the dynamics of its automorphism group.
7.1.1. Visual points
Let $\Omega\subset\mathbf{Sb}(G)$ be a proper domain. A point $p\in\partial\Omega$ will be called a visual point if it satisfies the following condition: for any pair of sequences $(x_{k}),(y_{k})$ of $\Omega$ such that $(x_{k})$ converges to $p$, if the sequence $(K_{\Omega}(x_{k},y_{k}))$ is bounded in $\mathbb{R}_{\geq 0}$, then $y_{k}\rightarrow p$.
Visual points satisfy the following geometric property:
Lemma 7.1.
Assume that $\Omega$ is a proper quasi-homogeneous domain of $\mathbf{Sb}(G)$. Let $p\in\partial\Omega$ be a visual point. Then $\operatorname{Z}_{p}\cap\Omega=\emptyset$.
  • Proof.
    Since $\Omega$ is quasi-homogeneous we can find $x\in\Omega$ and a sequence $(x_{k})\in\Omega^{\mathbf{N}}$ converging to $x$ such that, for some sequence $(g_{k})\in\mathsf{Aut}(\Omega)^{\mathbb{N}}$,
    $g_{k}\cdot x_{k}\underset{k\rightarrow+\infty}{\longrightarrow}p.$
    Since $p$ is visual, one has $g_{k}\cdot x\rightarrow p$. By invariance of $K_{\Omega}$ and the fact that $p$ is visual, one actually has $g_{k}\cdot y\rightarrow p$ for all $y\in\Omega$.
    Let $B\subset\Omega$ be the closed ball of center $x_{0}\in\Omega$ and of radius $1$ for the metric $K_{\Omega}$. If the sequence $g_{k}\cdot B$ does not converge to $\{p\}$ for the Hausdorff topology, then there is a neighborhood $\mathscr{V}$ of $p$ in $\mathbf{Sb}(G)$ and a sequence $(y_{k})\in B^{\mathbb{N}}$ such that $g_{k}\cdot y_{k}\notin\mathscr{V}$ for all $k\in\mathbb{N}$. Since $B$ is a compact subset of $\Omega$ by Corollary 6.13, up to extracting we may assume that there exists $y\in B$ such that $g_{k}\cdot y_{k}\rightarrow y$. Then $(g_{k}\cdot y)$ converges to $p$. But $K_{\Omega}(g_{k}\cdot y_{k},g_{k}\cdot y)\rightarrow 0$, and in particular the sequence $\left(K_{\Omega}(g_{k}\cdot y_{k},g_{k}\cdot y)\right)$ is bounded, so by visuality $g_{k}\cdot y_{k}\rightarrow p$, contradiction.
    Hence $g_{k}\cdot B\rightarrow\{x\}$ for the Hausdorff topology. Then, by the KAK decomposition, up to extracting, we may assume that there exists $q\in\mathbf{Sb}(G)$ such that $(g_{k})$ is $(p,q)$-contracting, in the sense that
    $g_{k|\mathbf{Sb}(G)\smallsetminus\operatorname{Z}_{q}}\rightarrow p,$
[…]
Remark 7.2.
(1) Lemma 7.1 is specific to our context: given a point $p$ of an arbitrary flag manifold $\mathscr{F}_{\Theta}$, the set $\operatorname{Z}_{p}$ is a subset of $\mathscr{F}_{\Theta}^{-}$, which cannot be $G$-equivariantly identified with $\mathscr{F}_{\Theta}$ in general. There are versions of Lemma 7.1 for flag manifolds that are not self-opposite, but they express a weaker geometric property for visual points (see e.g. [Gal25]). For instance, for the flag manifold $\mathbb{P}(\mathbb{R}^{n})$ with $n\geq 2$, which is as far as possible from being self-opposite, this weaker version of Lemma 7.1 expresses that any visual point of $\partial\Omega$ is contained in a supporting projective hyperplane to $\Omega$, which is already a consequence of the convexity of $\Omega$.
(2) The name “visual” refers to the visual boundary of the abstract metric space $(\Omega,K_{\Omega})$, defined as the set of all unbounded sequences of $(\Omega,K_{\Omega})$, quotiented by the relation
$(x_{k})\ \mathscr{R}\ (y_{k})\Longleftrightarrow(K_{\Omega}(x_{k},y_{k}))\text{ is bounded}.$
There is a natural embedding from the set of visual points of $\partial\Omega$ to the visual boundary of $(\Omega,K_{\Omega})$.
7.1.2. $\mathscr{R}$-extremal points
In [LZ19], a notion of $\mathscr{R}$-extremal point is defined using rank-one lines in the Grassmannians. A similar notion is introduced in [CG24], using the classical photons of the Einstein Universe. Here we define the analogous notion in the context of Shilov boundaries of $\operatorname{HTT}$ Lie groups:
Definition 7.3.
We say that a point $p\in\partial\Omega$ is $\mathscr{R}$-extremal if for any photon $\Lambda$ through $p$, the relative interior of $\Lambda\cap\partial\Omega$ in $\Lambda$ does not contain $p$.
Following the notation of [LZ19], we denote by $\operatorname{Extr}_{\mathscr{R}}(\Omega)$ the set of $\mathscr{R}$-extremal points of $\Omega$. We will see in Section 7.2 that this set is never empty. In this section we aim at showing the following link between visuality and $\mathscr{R}$-extremality:
Proposition 7.4.
Let $\Omega\subset\mathbf{Sb}(G)$ be a proper quasi-homogeneous domain. Then any $\mathscr{R}$-extremal point of $\Omega$ is visual.
Remark 7.5.
  1. (1)
    Theorem 1.4 will actually imply that the reciprocal of Proposition 7.4 is true.
  2. (2)
    Proposition 7.4 is an analogue of classical Vey’s Lemma (Fact 1.8) of real projective geometry, in the context of Shilov boundaries of Hermitian symmetric spaces. Note that in Proposition 7.4, we assume that $\Omega$ is quasi-homogeneous, which is not needed in Vey’s lemma. The reason lies in the definition of the Kobayashi metric: given two sequences $(x_{n}),(y_{n})$ of $\Omega$ such that $x_{n}\rightarrow p\in\operatorname{Extr}_{\mathscr{R}}(\Omega)$, the minimal integer $N\in\mathbb{N}$ such that $\mathscr{C}_{x_{n},y_{n}}^{N}(\Omega)$ is nonempty could go to infinity while the distance $K_{\Omega}(x_{n},y_{n})$ remains bounded. Then it could be possible for $(y_{n})$ to converge to a point $q\in\partial\Omega\smallsetminus\{p\}$. This phenomenon is avoided when $\Omega$ is quasi-homogeneous.
For the proof of Proposition 7.4, we follow the strategy of [LZ19, Thm 7.4]. We will need the following definition:
Definition 7.6.
Let $\Omega$ be a proper domain of $\mathbf{Sb}(G)$ and $x,y\in\Omega$, and $N\in\mathbb{N}^{*}$. Let us define
$K_{\Omega}^{N}(x,y):=\inf\left\{\mathsf{len}_{\Omega}(\gamma)\big{|}\,\gamma\in\mathscr{C}_{x,y}^{N}(\Omega)\right\}.$
Recall that $\mathsf{len}_{\Omega}(\gamma)$ is the $K_{\Omega}$-length for the metric $K_{\Omega}$ of the path $\gamma$ (see Section 6.1.4). The quantity $K_{\Omega}^{N}(x,y)$ is finite if, and only if, the set $\mathscr{C}_{x,y}^{N}(\Omega)$ is nonempty.
The map $K_{\Omega}^{N}:\Omega\times\Omega\rightarrow\mathbb{R}\cup\{\infty\}$ is $\mathsf{Aut}(\Omega)$-invariant. The sequence $(K_{\Omega}^{N}(x,y))_{N\in\mathbb{N}}$ is nonincreasing, eventually finite, and one has $K_{\Omega}(x,y)=\lim_{N\rightarrow+\infty}K_{\Omega}^{N}(x,y)$.
The first step in the proof of Proposition 7.4 is a weaker version, where we replace $K_{\Omega}$ by $K^{N}_{\Omega}$ (note that on the other hand we do not need the quasi-homogeneity assumption in the next lemma):
Lemma 7.7.
Let $\Omega\subset\mathbf{Sb}(G)$ be a proper dually convex domain. Let $p\in\partial\Omega$ be an $\mathscr{R}$-extremal point, and let $q\in\overline{\Omega}$. Let $(x_{k}),(y_{k})\in\Omega^{\mathbb{N}}$ be two sequences such that $x_{k}\rightarrow p$ and $y_{k}\rightarrow q$, and such that there exist $N\in\mathbb{N}$ and $M>0$ such that $K_{\Omega}^{N}(x_{k},y_{k})\leq M$ for all $k\in\mathbb{N}$. Then $p=q$.
  • Proof.
    For any $k\in\mathbb{N}$, let $\gamma_{k}=(x_{k}^{0}:=x_{k},\dots,x_{k}^{N}:=y_{k})\in\mathscr{C}^{N}_{x_{k},y_{k}}(\Omega)$ be such that
    $\sum_{i=0}^{N-1}K_{\Omega}(x_{i}^{k},x_{i+1}^{k})=\mathsf{len}_{\Omega}(\gamma_{k})\leq K_{\Omega}^{N}(x_{k},y_{k})+1\leq M+1,$
    the first equality holding because of (6.8). Then, one has $K_{\Omega}(x_{k}^{i},x_{k}^{i+1})\leq M+1$ for all $10\leq i\leq N-$ 1. Hence one can assume that $N=1$, and the lemma follows by induction.
    Let us then assume that $N=1$. For all $k$, the two points $x_{k}$ and $y_{k}$ lie in the same connected component of the intersection $I_{k}:=\Lambda_{k}\cap\Omega$ of a photon $\Lambda_{k}$ with $\Omega$. Let $a_{k},b_{k}$ be the endpoints of $I_{k}$ such that $a_{k},x_{k},y_{k},b_{k}$ are aligned in this order. If $g_{k}\in G$ is such that $\Lambda_{k}=g_{k}\cdot\Lambda_{\mathsf{std}}$ and if we define $\mathsf{i}^{k}:=\mathsf{i}_{g_{k}^{-1}}$ (recall (5.1)), then there exist $r_{k},s_{k},t_{k},u_{k}\in\mathbb{P}(\mathbb{R}^{2})$, aligned in this order, such that
    $\mathsf{i}^{k}(r_{k})=a_{k};\quad\mathsf{i}^{k}(s_{k})=x_{k};\quad\mathsf{i}^{k}(t_{k})=y_{k};\quad\mathsf{i}^{k}(u_{k})=b_{k}.$
    Then Proposition 6.8 implies that
    $\log(r_{k}:s_{k}:t_{k}:u_{k})=\mathsf{k}_{\Omega}(x_{k},y_{k})=K_{\Omega}(x_{k},y_{k})\leq M.$
[…]
To prove Proposition 7.4, we need to strengthen Lemma 7.7 and replace $K_{\Omega}^{N}$ by $K_{\Omega}$ in the assumptions. Lemma 7.9 below, together with the quasi-homogeneity assumtion, will allow us to do so. To prove Lemma 7.9, we prove the auxiliary Lemma 7.8.
Since $K_{\Omega}$ is continuous, one can define the diameter of a compact subset $\mathscr{C}$ of $\Omega$ as:
$\operatorname{diam}(\mathscr{C}):=\max\left\{K_{\Omega}(x,y)\mid x,y\in\mathscr{C}\right\}.$
Then one has:
Lemma 7.8.
Let $\Omega\subset\mathbf{Sb}(G)$ be a proper dually convex domain, and let $D$ be a diamond such that $\overline{D}\subset\Omega$. Then for any $x,y\in D$, there exists an $\mathsf{n}(G)$-chain from $x$ to $y$ with $K_{\Omega}$-length less than or equal to $\mathsf{n}(G)\operatorname{diam}(\overline{D})$. In particular,
$K_{\Omega}^{\mathsf{n}(G)}(x,y)\leq\mathsf{n}(G)\operatorname{diam}(\overline{D}).$
  • Proof.
    We know by Lemma 6.2 that there exists an $\mathsf{n}(G)$-chain $\gamma=(x_{0},\cdots,x_{N})$ from $x$ to $y$ contained in $D$. By Equation Proposition 6.8, one has
    $K_{\Omega}^{\mathsf{n}(G)}(x,y)\leq\mathsf{len}_{\Omega}(\gamma)=\sum_{i=0}^{N-1}K_{\Omega}(x_{i},x_{i+1})\leq\sum_{i=0}^{N-1}K_{D}(x_{i},x_{i+1})\leq\mathsf{n}(G)\operatorname{diam}(\overline{D}).$
[…]
Lemma 7.9.
Let $\Omega\subset\mathbf{Sb}(G)$ be a proper dually convex domain, and let $\mathscr{K}\subset\Omega$ be a compact subset. Then there exist $N\in\mathbb{N}$ and $M>0$ such that for any $x,y\in\mathscr{K}$,
$K_{\Omega}^{N}(x,y)\leq M.$
  • Proof.
    First choose, for any $x\in\mathscr{K}$, a diamond $D_{x}$ of diameter less than 1 for $K_{\Omega}$, and containing $x$. Such a diamond exists by Remark 4.9.(1) and Proposition 6.7.
    For any $x,y\in\mathscr{K}$, there is a minimal $N(x,y)\in\mathbb{N}$ such that
    $M(x,y):=K_{\Omega}^{N(x,y)}(x,y)<+\infty.$
    Let $(z,z^{\prime})\in D_{x}\times D_{y}$. There exists $\gamma\in\mathscr{C}_{x,y}^{N(x,y)}(\Omega)$ such that $\mathsf{len}_{\Omega}(\gamma)\leq M(x,y)+1$. Besides, by Lemma 7.8, there exist $\gamma_{1}\in\mathscr{C}_{z,x}^{\mathsf{n}(G)}(\Omega)$ and $\gamma_{2}\in\mathscr{C}_{y,z}^{\mathsf{n}(G)}(\Omega)$, both of $K_{\Omega}$-length less than or equal to $\mathsf{n}(G)$. If $N_{0}:=N(x,y)+2\mathsf{n}(G)$ and $\gamma_{3}$ is the concatenation of $\gamma_{1}$, $\gamma$ and $\gamma_{2}$, then $\gamma_{3}\in\mathscr{C}_{z,z^{\prime}}^{N_{0}}(\Omega)$ and for any $N\geq N$ one has
    $K^{N}_{\Omega}(z,z^{\prime})\leq K^{N_{0}}_{\Omega}(z,z^{\prime})\leq\mathsf{len}_{\Omega}(\gamma_{3})=\mathsf{len}_{\Omega}(\gamma_{1})+\mathsf{len}_{\Omega}(\gamma)+\mathsf{len}_{\Omega}(\gamma_{2})\leq 2\mathsf{n}(G)+M(x,y)+1.$
    Let $F\subset\mathscr{K}$ be a finite subset such that $\mathscr{K}\subset\bigcup_{x\in F}D_{x}$. Then for all $z,z^{\prime}\in\mathscr{K}$, one has $K_{\Omega}^{N}(z,z^{\prime})\leq M$, where
    $N$ $:=\max\{N(x,y)\mid x,y\in F\}+2\mathsf{n}(G);$
    $M$ $:=\max\{M(x,y)\mid x,y\in F\}+2\mathsf{n}(G)+1.\quad\hfill\square$
  • Proof Proposition 7.4.
    Note that $\Omega$ is proper and dually convex by Proposition 4.4, hence Lemmas 7.7 and 7.9 apply to $\Omega$. Write $\Omega=\mathsf{Aut}(\Omega)\cdot\mathscr{K}$ where $\mathscr{K}\subset\Omega$ is a compact subset. Let $(x_{k})\in\Omega^{\mathbb{N}}$ be such that $x_{k}\rightarrow p$, and let $(y_{k})\in\Omega^{\mathbb{N}}$ be such that there exists $M>0$ with $K_{\Omega}(x_{k},y_{k})\leq M$ for all $k\in\mathbb{N}$. For all $k\in\mathbb{N}$, there exists $g_{k}\in\mathsf{Aut}(\Omega)$ such that $g_{k}\cdot x_{k}\in\mathscr{K}$. Let
    $\mathscr{K}^{\prime}:=\{x\in\Omega\mid K_{\Omega}(x,\mathscr{K})\leq M\}.$
    The set $\mathscr{K}^{\prime}$ is compact by properness of $K_{\Omega}$ (see Corollary 6.13). Then, by Lemma 7.9, there exist $N\in\mathbb{N}$ and $M^{\prime}>0$ such that $K^{N}_{\Omega}(a,b)\leq M^{\prime}$ for all $a,b\in\mathscr{K}^{\prime}$. By $\mathsf{Aut}(\Omega)$-invariance of $K_{\Omega}$, for any $k\in\mathbb{N}$ one has $g_{k}\cdot x_{k},\ g_{k}\cdot y_{k}\in\mathscr{K}^{\prime}$. Hence by $\mathsf{Aut}(\Omega)$-invariance of $K_{\Omega}^{N}$ one has
    $K_{\Omega}^{N}(x_{k},y_{k})=K_{\Omega}^{N}(g_{k}\cdot x_{k},\ g_{k}\cdot y_{k})\leq M^{\prime}.$
[…]
7.2. A special set of $\mathscr{R}$-extremal points: the strongly $\mathscr{R}$-extremal points
Let $\Omega\subset\mathbf{Sb}(G)$ be a domain which is proper in $\mathbb{A}$. We say that a point $p\in\partial\Omega$ is strongly $\mathscr{R}$-extremal if either $\mathbf{C}^{-}(p)\cap\overline{\Omega}=\{p\}$ or $\mathbf{C}^{+}(p)\cap\overline{\Omega}=\{p\}$.
In general there are less strongly $\mathscr{R}$-extremal points than $\mathscr{R}$-extremal points. However, the next lemma shows that strongly $\mathscr{R}$-extremal points always exist.
Lemma 7.10.
Let $\Omega$ be a domain which is proper in $\mathbb{A}$. Then for any $x\in\Omega$ there exist two strongly $\mathscr{R}$-extremal points $p\in\mathbf{J}^{-}(x)$ and $q\in\mathbf{J}^{+}(x)$.
  • Proof.
    Up to translating $\Omega$ in $\mathbb{A}$, we may assume that $x=P$. Since $c^{0}$ is a properly convex cone of $\mathfrak{u}^{-}$, there is a nonzero linear map $\psi$ of $\mathfrak{u}^{-}$ such that $\overline{c^{0}}\smallsetminus\{0\}$ is contained in $\{\psi>0\}$. Let $X\in\mathfrak{u}^{-}$ be the element of $\overline{\varphi_{\mathsf{std}}^{-1}(\Omega)}\cap(-\overline{c^{0}})$ such that $\psi(X)$ is minimal. Then $p:=\exp(X)P$ lies in $\partial\Omega$.
    Let us show that $p$ is strongly $\mathscr{R}$-extremal. Let $y\in\mathbf{C}^{-}(p)$. Write $y=\exp(Y)P$ with $Y\in X-\partial c^{0}$. Then one has
    (7.1) $\psi(X-Y)\geq 0,\text{ with equality if, and only if, }y=p.$
[…]
Remark 7.11.
  1. (1)
    Lemma 7.10 is a generalization of [CG24, Lem. 6.4], where the $\mathscr{R}$-extremal points are called photon-extremal.
  2. (2)
    By Lemma 5.5, Remark 5.6 and the fact that $c^{0}$ is a properly convex cone in $\mathfrak{u}^{-}$, strongly $\mathscr{R}$-extremal points are always $\mathscr{R}$-extremal, but the converse is false in general. For instance, for $G=\operatorname{SO}(n,2)$, take $p,q\in\mathbb{A}$ with $q\in\mathbf{I}^{+}(p)$. Then $\mathbf{D}(p,q)$ has exactly two strongly $\mathscr{R}$-extremal points, namely $p$ and $q$. The points of $\mathbf{C}^{+}(p)\cap\mathbf{C}^{-}(q)$ are $\mathscr{R}$-extremal but not strongly $\mathscr{R}$-extremal.
  3. (3)
    Contrary to the notion of $\mathscr{R}$-extremality, that of strong $\mathscr{R}$-extremality is only defined for a domain $\Omega$ which is proper in $\mathbb{A}$. It is not clear at first that this second notion is invariant under $\mathsf{Aut}(\Omega)$. Using Lemma 3.5 it is actually possible to show that this is the case whenever $\Omega$ is dually convex, although we will not use this fact. We will only prove this invariance in the quasi-homogeneous case; see Lemma 8.2.
8. End of the proof of Theorem 1.4
In this section we finish the proof of Theorem 1.4. We take $\Omega\subset\mathbf{Sb}(G)$ a quasi-homogeneous domain, proper in $\mathbb{A}$.
Let $x\in\Omega$ and let $p_{0}\in\partial\Omega\cap\mathbf{J}^{-}(x)$ and $q_{0}\in\partial\Omega\cap\mathbf{J}^{+}(x)$ be two strongly $\mathscr{R}$-extremal points of $\partial\Omega$ given by Lemma 7.10. Then in particular $p_{0},q_{0}\in\operatorname{Extr}_{\mathscr{R}}(\Omega)$ (Remark 7.11.(2)), so by Lemma 7.1 and Proposition 7.4, one has
(8.1) $\Omega\cap\operatorname{Z}_{p_{0}}=\Omega\cap\operatorname{Z}_{q_{0}}=\emptyset.$
By reflexivity, one has $x\in\mathbf{J}^{+}(p_{0})$. By (8.1), we also know that $x\notin\operatorname{Z}_{p_{0}}$. Then, by Fact 3.4, one has $x\notin\mathbf{C}(p_{0})$, and hence $x\in\mathbf{I}^{+}(p_{0})$. Similarly, one has $x\in\mathbf{I}^{-}(q_{0})$. Hence $x\in\mathbf{D}(p_{0},q_{0})$. By connectedness of $\Omega$, we then have the inclusion
(8.2) $\Omega\subset\mathbf{D}(p_{0},q_{0}).$
The goal of the rest of this section is to prove the converse inclusion. First observe that $p_{0}$ and $q_{0}$ are characterized among $\mathscr{R}$-extremal points of $\partial\Omega$ by a geometric property:
Lemma 8.1.
Let $p\in\operatorname{Extr}_{\mathscr{R}}(\Omega)$ be such that $\mathbf{I}^{+}(p)\cap\Omega\neq\emptyset$ (resp. $\mathbf{I}^{-}(p)\cap\Omega\neq\emptyset$). Then $p=p_{0}$ (resp. $p=q_{0}$).
  • Proof.
[…]
Using Lemma 3.5, we can now prove:
Proposition 8.2.
Let $g\in\mathsf{Aut}(\Omega)$. Then $g\cdot p_{0},g\cdot q_{0}\in\{p_{0},q_{0}\}$.
  • Proof.
[…]
We have shown in Proposition 8.2 that for any $g\in\mathsf{Aut}(\Omega)$, the element $g$ stabilizes the pair $\{p_{0},q_{0}\}$. Then $\mathsf{Aut}(\Omega)$ preserves the set $\mathbf{Sb}(G)\smallsetminus(\operatorname{Z}_{p_{0}}\cup\operatorname{Z}_{q_{0}})$, and hence permutes its connected component. Since $\Omega\subset\mathbf{D}(p_{0},q_{0})$ is $\mathsf{Aut}(\Omega)$-invariant, the group $\mathsf{Aut}(\Omega)$ preserves the connected component $\mathbf{D}(p_{0},q_{0})$ of $\mathbf{Sb}(G)\smallsetminus(\operatorname{Z}_{p_{0}}\cup\operatorname{Z}_{q_{0}})$. Then:
(8.3) $\mathsf{Aut}(\Omega)\leq\mathsf{Aut}(\mathbf{D}(p_{0},q_{0})).$
The next proposition ends the proof of Theorem 1.4.
Proposition 8.3.
We have $\Omega=\mathbf{D}(p_{0},q_{0})$.
  • Proof.
    Since $\Omega$ is open in the domain $D:=\mathbf{D}(p_{0},q_{0})$, it suffices to prove that $\Omega$ is closed in $D$. Let $(x_{k})\in\Omega^{\mathbb{N}}$ converging to a point $x\in D$. Write $\Omega=\mathsf{Aut}(\Omega)\cdot\mathscr{K}$, with $\mathscr{K}\subset\Omega$ a compact subset. For any $k\in\mathbb{N}$ there exist $(g_{k},z_{k})\in\mathsf{Aut}(\Omega)\times\mathscr{K}$ such that $g_{k}\cdot z_{k}\rightarrow x$. Up to extracting there exists $z\in\mathscr{K}$ such that $z_{k}\rightarrow z$. Then
    $K_{D}(g_{k}\cdot z,x)\leq K_{D}(g_{k}\cdot z,g_{k}\cdot z_{k})+K_{D}(g_{k}\cdot z_{k},x)=K_{D}(z,z_{k})+K_{D}(g_{k}\cdot z_{k},x)\underset{k\rightarrow+\infty}{\longrightarrow}0.$
    The last equality holds by Proposition 6.5.(2) and (8.3), and the limit by Proposition 6.7. Since $K_{D}$ generates the standard topology on $D$ (again by Proposition 6.7), we have $g_{k}\cdot z\rightarrow x$.
[…]
Remark 8.4.
The arguments of this section generalize those of [CG24]. However, the causal convexity arguments used in the latter, and specifically [CG24, Prop. 6.2], do not hold anymore. Instead, in the present paper, causality in $\mathbf{Sb}(G)$ appears in the proofs of Propositions 8.2 and 8.3.
9. Application: closed manifolds with proper development
In this section we prove Corollary 1.7. A manifold $M$ is a $(G,\mathbf{Sb}(G))$-manifold, if there exists a (maximal) atlas of charts $(U,\psi_{U})_{U\in\mathscr{A}}$ on $M$ with values in $\mathbf{Sb}(G)$, such that for any $U,V\in\mathscr{A}$ with $U\cap V\neq\emptyset$, the map $\psi_{V}\circ\psi_{U}^{-1}$ is the restriction of an element of $G$ to $\varphi_{U}(U\cap V)$. In this case there exists a map $\mathsf{dev}:\widetilde{M}\rightarrow\mathbf{Sb}(G)$, called the developing map (unique up to postcomposition by elements of $G$), where $\widetilde{M}$ is the universal cover of $M$. We say that $M$ is proper if $\mathsf{dev}(\widetilde{M})$ is a proper domain of $\mathbf{Sb}(G)$. This property is independent of the the choice of the developing map for $M$.
  • Proof of Corollary 1.7.
    The corollary is straightforward if $G$ is of real rank 1, hence we may assume that the real rank of $G$ is $r\geq 2$.
    Let $\pi_{1}(M)$ be the fundamental group of $M$ and let $\mathsf{hol}:\pi_{1}(M)\rightarrow G$ be the holonomy of $M$, so that the map $\mathsf{dev}$ is $\mathsf{hol}$-equivariant. Let $\Omega=\mathsf{dev}(\widetilde{M})$. Since $M$ is closed, there is a compact fundamental domain $\mathscr{K}\subset\widetilde{M}$ intersecting every $\pi_{1}(M)$-orbit. By $\mathsf{hol}$-equivariance of $\mathsf{dev}$, the compact set $\mathsf{dev}(\mathscr{K})\subset\Omega$ intersects any $\mathsf{hol}(\pi_{1}(M))$-orbit. Hence $\Omega$ is quasi-homogeneous. Theorem 1.4 implies that $\Omega$ is a diamond and is $\left(H_{1},H_{2}\right)$-equivariantly diffeomorphic to the symmetric space ×$\mathbb{R}\times X_{L_{s}}$, where $H_{1},H_{2}$ are finite-index subgroupds of $\mathsf{Aut}(\Omega),\mathrm{Isom}\left(\mathbb{R}\times X_{L_{s}}\right)$ respectively. Up to taking a finite cover of $M$, we may assume that $H_{1}=\mathsf{Aut}(\Omega)$ and $H_{2}=\mathrm{Isom}(\mathbb{R}\times X_{L_{s}})$.
    Let $g_{\Omega}$ be the invariant Riemannian metric of $\Omega$ equal to $g_{\mathbb{R}}\oplus g_{X_{L_{s}}}$ under the previous identification and let $g=\mathsf{dev}^{*}g_{\Omega}$. The metric $g$ is invariant under $\pi_{1}(M)$, so it defines a Riemannian metric on $M$. This metric must be complete since $M$ is closed, so $g$ is also complete. The map $\mathsf{dev}$ is a local isometry between complete Riemannian manifolds, so it is a covering map. Since $\Omega$ is simply connected, the covering map is a diffeomorphism onto its image and $M$ is a quotient of $\Omega$ by a cocompact lattice of $\mathsf{Aut}(\Omega)$.
[…]